### v1

# Microbenchmarking NVIDIA’s Blackwell Architecture: An in-depth Architectural Analysis

Abstract
As GPU architectures rapidly evolve to meet the overcoming demands of exascale computing and machine learning, the performance implications of architectural innovations remain poorly understood across diverse workloads. NVIDIA’s Blackwell (B200) generation introduce significant architectural advances including the 5th generation tensor cores, tensor memory (TMEM), decompression engine (DE), and dual chips; however systematic methodologies for quantifying these improvements lag behind hardware development cycles. We contribute an open-source microbenchmark suite that offers practical insights into optimizing workloads to fully utilize the rich feature sets of the modern GPU architecture. This work aims to enable application developers make informed architectural decisions and guide future GPU design directions.
Our work studies Blackwell GPUs, compares them to H200 generation with regards to the memory subsystem, tensor core pipeline and floating-point precisions (FP32, FP16, FP8, FP6, FP4). Our systematic evaluation of dense/sparse GEMM, transformer inference, and training workloads demonstrate that B200’s tensor core enhancements achieves 1.56× higher mixed-precision throughput and 42% better energy efficiency than H200. Our memory analysis reveals 58% reduction in memory access latency in cache-misses, fundamentally changing optimal algorithm design strategies.
I Introduction
Artificial Intelligence (AI) and high‑performance computing (HPC) have evolved into data‑intensive disciplines that continuously challenge hardware efficiency, scalability, and precision. Large language models (LLMs) now exceed hundreds of billions of parameters and process context windows spanning millions of tokens [ding2024longrope, kevian2024capabilitieslargelanguagemodels], alongside multi‑physics and climate simulations that demand teraflops of sustained performance, thus shifting GPU design to enable both massive parallel and architectural adaptability. At these scales, modern accelerators must balance several demands: maintaining arithmetic throughput for dense tensor workloads, minimizing on‑chip and off‑chip memory latency, while offering hardware primitives that effectively support mixed‑precision computation.
The growing demands have exposed several limitations of current GPU architectures, particularly within their memory hierarchies, precision flexibility, and latency‑sensitive task scheduling. As a result, sustained architectural innovation in accelerators has become essential for advancing both throughput‑optimized training and time‑critical inference workloads. One such architecture that is designed to address some of these challenges is NVIDIA’s Blackwell architecture [NVIDIA2024_Blackwell] that showcases a major generational evolution.
As the direct successor to the Hopper generation, the Blackwell architecture extends NVIDIA’s GPU design in several modifications across the compute pipeline, memory hierarchy, and tensor processing subsystems. Blackwell introduces 5th‑generation Tensor Cores capable of FP4 and FP6 precision execution, offering trade‑offs between accuracy and performance for large‑scale training. In addition, the introduction of the Tensor Memory (TMEM) subsystem as a dedicated on‑chip memory for tensor data movement reduces reliance on shared memory (SMEM) and per SM register files (RF) during matrix‑intensive operations. Next, NVIDIA included a hardware decompression engine (DE) and redesigned the instruction pipeline for access to compressed model weights. Beyond raw compute enhancements, Blackwell also has a revised thread and CTA scheduling model to utilize inter‑SM communication and memory concurrency. With so many changes introduced, intended to address the escalating demands of AI, gaming and scientific computing, an analysis of the microarchitecture and new instructions is necessary, which will provide application developers and scientists to achieve the highest performance possible for modern and future GPUs.
This paper introduces a newly developed open-source microbenchmark suite (unable to share the code at this time due to double-blind), implemented in PTX and CUDA, that enables comprehensive architectural analysis of NVIDIA’s Blackwell GPU. Emphasizing innovations that distinguish it from Hopper, the suite systematically evaluates performance under stress—particularly in compute-bound and memory-bound workloads—revealing implications for parallel computing applications.
The key contributions of our work are as follows:
Build targeted microbenchmarks to characterize key components of NVIDIA Blackwell B200, to the best of our knowledge - our work is the first detailed microbenchmark characterization of this next-generation GPU.
Quantify TMEM’s impact on matrix-heavy workloads and its role in reducing memory bottlenecks in tensor computations.
Evaluate the decompression engine’s throughput across formats and identify optimal usage.
Analyze 5th-gen Tensor Core execution via the new tcgen05 PTX instructions to study performance implications
Assess FP4/FP6 performance and accuracy trade-offs in mixed-precision tensor operations quantifying accuracy-performance trade-offs
Benchmark Blackwell across LLM inference/training, scientific kernels, and mixed-precision workloads to demonstrate real-world impact and performance gains.
Provide actionable performance guidelines for developers leveraging Blackwell’s architecture.
The remainder of this paper is organized as follows: Section II details our contributions to the current state of the art GPU microbenchmarks. Following, Section III presents an overview of the Blackwell’s B200 architecture. After which, Section IV details the microbenchmark methodology we employ to systematically characterize the Blackwell microarchitecture. While, Section V details the memory subsystem before Section VI presents the Tensor Core pipeline. Section VII presents performance analysis across key workloads and to conclude we discuss implications and trade-offs in Section VIII.
II Related Work
Understanding GPU performance has long been a critical focus in HPC research. Over the years, several studies have used microbenchmarks and other methodologies to dissect architectural layers and analyze GPU microarchitectures in fine-grained detail. Early studies on Tesla and Fermi focused on memory and cache behavior [Wong2010, Subramoniapillai2012], while later work dissecting Kepler, Pascal [Zhang2017], and Maxwell [Mei2012] examined warp scheduling and instruction latency. With Turing through Hopper [Fasi2021, jia2019, tan2011, Markidis2018, Martineau2018, Raihan2019, Yan2020, luo2025], research shifted to mixed-precision and tensor core performance, introducing benchmarks for mma instructions, tile sizes, and data layouts. Recent efforts also explore instruction-level parallelism[Sun2022], and pipeline dynamics under high register pressure.
Beyond microbenchmarking, researchers built frameworks to characterize GPU performance. Application profiling[5336219] gathers runtime metrics but faces overhead and limited architectural visibility. Roofline models [leinhauser2021metricsdesigninstructionroofline] offer throughput vs. intensity plots, yet oversimplify bottlenecks and miss dynamic memory behaviors. Cache stall prediction [wenhao2012] estimates pipeline delays from access patterns but fails to capture modern GPU complexities like cache bypassing, warp scheduling, and memory coalescing.
Analytical models like Accel-Sim [accelsim2020] and GCoM [gcom2022], built on Hong and Kim’s work [Hong2009], offer useful GPU performance insights but neither of them model Blackwell-specific features like TMEM or the DE, as the detailed architectural information required for accurate simulation remain unknown.
Thus, without a systematic understanding of these components, the research community lacks critical data needed for performance modeling, workload optimization, and accurate simulation of AI reasoning workloads typically needed for datacenter deployments.
III Blackwell Architecture
In this Section, we introduce the architecture of the data center NVIDIA B200 GPU, based on the NVIDIA Blackwell Architecture, and then detail the divergence from prior designs.
III-A Blackwell Architecture
The B200 GPU signifies a decisive progression in architectural philosophy. Previously, GPU generations from Tesla to Hopper focused on maximizing floating-point operations per second (FLOPS) for large-scale model training. In contrast, Blackwell emphasizes post-training and inference efficiency, adopting transformational changes in both memory and compute organization. One B200 GPU includes a dual-die configuration [NVIDIA2024_Blackwell] where two GPU dies comprise 208 billion transistors, feature 148 SMs spread across eight GPCs, provide four L2 cache partitions (double those in Hopper), and include eight HBM3e memory stacks. Though physically partitioned, both dies are unified by the NVIDIA High-Bandwidth Interface (NV-HBI) providing a coherent and single device to software, with unified 192 GB HBM3e memory space.
Within each SM, Blackwell introduces fifth-generation Tensor Cores that break from the warp-synchronous paradigm characterizing earlier architectures (Volta, Ampere, Hopper). Previous generations enforced that all 32 threads within a warp synchronize before executing matrix multiply–accumulate (MMA) operations via the mma.sync or wgmma instructions. This lock-step model reduced scheduling flexibility and created idle cycles, especially for dependency chains of varying lengths.
Blackwell replaces warp-synchronous MMA with tcgen05.mma, a single-thread instruction. Now, each thread independently issues MMA operations, removing warp-level synchronization and enabling true per-thread scheduling for tensor operations. Operands are now supplied from shared memory (SMEM) and a new memory pathway: Tensor Memory (TMEM). Per SM, the TMEM provides memory access to and from tensor cores. Allocation, data movement, and deallocation are explicitly managed in software via the tcgen PTX set of instructions, giving compiler toolchains precise control over tile locality and traffic patterns.
The flexibility of independent MMA dispatch reduces idle cycles and exposes optimization opportunities for the compiler, though it also raises questions on new performance limits: instruction latency under dependency, concurrency of tensor core usage, and pipeline saturation. These remain undocumented in vendor literature and are explored in our systematic characterization.
In terms of numerical support, Blackwell’s Tensor Cores introduce native 4-bit and 6-bit floating-point precision (FP4 and FP6) for quantized inference, further improving memory and computational efficiency for AI workloads. Architectural innovations extend to the thread-block level, with CTA pair execution: two Cooperative Thread Arrays (CTAs) with adjacent ranks share operands, reducing redundant data movement. Each CTA pair maps to a TPC and leverages a dedicated intra-TPC communication network for efficient operand sharing.
Further broadening the functionality, Blackwell’s Tensor Cores provide native support for convolution operators with weight-stationary dataflows that use a collector buffer to cache and reuse matrix B (weight tensor) operands. Hence, optimizing for convolutional kernels that benefit from operand locality. Blackwell also addresses growing model and data sizes by introducing a hardware-based Decompression Engine (DE) to offload decompression tasks from general-purpose SMs. This subsystem supports various algorithms, more on that in Section V, enabling model weights and large database tables to be stored in compressed form within HBM3e and decompressed transparently during memory access [NVIDIA2024_Blackwell].
While some architectural details are publicly disclosed, critical microarchitectural information, such as instruction latency, pipeline depth, cache interaction, and saturation, remains unknown. Our PTX microbenchmark experiments (Sections V- VI) provide a systematic investigation to fill these knowledge gaps as they relate to AI and HPC performance.
IV PTX-Microbenchmark Methodology
We employ a microbenchmarking approach based on NVIDIA’s Parallel Thread Execution (PTX) to characterize Blackwell’s microarchitectural features. While prior GPU characterization studies [Wong2010, Mei2012, Yan2020] established foundational PTX-level benchmarking principles, we extend these methodologies with novel techniques specifically designed to dissect Blackwell’s previously uncharacterized components, including the 5th-generation Tensor Cores’s FP4/FP6 precision modes, the DE, and the revised cache hierarchy.
Our approach leverages PTX to provide explicit control over registers and memory operations specific to architectures. PTX code is compiled into Streaming Assembler (SASS) instructions or machine code executed by the GPU. We ensure our microbenchmarks accurately isolate and measure specific microarchitectural behaviors by documenting the PTX-to-SASS translation and validating against observed performance.
IV-A Novel Benchmark Design for Blackwell-Specific Features
IV-A1 Tensor Memory (TMEM)
Unlike previous architectures where MMA operations relied exclusively on SMEM, DSMEM, and RFs, Blackwell introduces TMEM as a dedicated on-chip memory specifically for tensor operations. This necessitates new data movement paradigms and presents unexplored opportunities for performance optimization. Understanding TMEM’s performance characteristics is critical for several reasons. First, traditional data movement instructions (including wmma.load, ldmatrix, ld.shared, and cp.async) which cannot interface with TMEM. Thus developers are required to adopt entirely new instruction sequences (tcgen05.ld, tcgen05.st, tcgen05.cp). Second, the performance implications of this new memory tier remain uncharacterized, leaving application developers without guidance on when and how to leverage TMEM effectively.
Our work addresses this gap by providing the first comprehensive empirical analysis of TMEM performance characteristics and their impact on real-world computational kernels.
Our methodology comprises of three evaluation strategies.
Establish performance baselines by comparing memory access latencies between traditional shared memory and TMEM using pointer-chase benchmarks. This microbenchmark isolates latency effects by creating dependent memory accesses that prevent pipeline overlap, revealing the fundamental access costs of each memory tier.
Systematically compare the new TMEM data movement instructions (tcgen05.* family) against their predecessors across varying access patterns.
With varying operand sizes and access strides, bandwidth saturation points are identified and per-access latency across different configurations is measured. This reveals both the capabilities and limitations of the new instruction set.
IV-A2 Decompression Engine Characterization
To systematically characterize the B200’s hardware DE, we develop a custom microbenchmarking suite targeting seven compression formats (LZ4, Snappy, Zstandard, GZIP, Cascaded, Bitcomp, ANS) across controlled test conditions. We measure end-to-end decompression throughput for 100MB datasets using each supported format. Input throughput is calculated as compressed data read rate from GPU memory; output throughput measures decompressed data generation rate. Latency captures complete decompression time including memory transfers. Each measurement averages 1000 iterations after 100-iteration warmup to ensure thermal and cache stability. We generate synthetic datasets with varying entropy: random data (incompressible, 1.00x ratio), mixed alphanumeric (1.98x), repetitive patterns (15.02x), and zero-filled buffers (245.45x).
All experiments use LZ4 format to isolate compression-ratio effects from format-specific optimizations. Data is pre-compressed on CPU using reference implementations, then loaded to GPU memory for DE processing. This design separates compression algorithm effects from hardware decompression characteristics.
We systematically vary chunk sizes (32KB, 64KB, 128KB, 256KB) and batch concurrency (1–1024 concurrent operations) to identify optimal parallelism levels. Peak throughput is measured at maximum sustainable bandwidth before efficiency degradation. Pipeline depth represents the concurrency level maintaining roughly 85% efficiency (defined as throughput per operation / peak single-operation throughput). Saturation point identifies where additional concurrency yields around 5% marginal throughput improvement. This methodology reveals hardware resource limits and memory bandwidth constraints not documented by NVIDIA.
IV-A3 Tensor Core Characterization
We develop custom GPU kernels to execute MMA operations of the form using Blackwell’s newly introduced tensor core instruction set (tcgen05).
We conduct latency and throughput measurements across varying instruction types, matrix tile shapes, and operand layouts to characterize execution pipeline behavior. Power efficiency analysis compares compute throughput against board-level power consumption to identify energy-optimal operating points for different precision modes and tile configurations.
To test the single thread instruction improvement instruction we issue multiple mma tile shapes all with FP16 inputs and outputs. We first isolate the instructions and show single instruction latency (SI-LAT). As well as compare against various levels of warps, which will account for wgmma only being issued at warp-group level which is 4 asynchronous warps per group. As well as using the TMEM as minimal as allowed, which means only for the accumulator, denoted as SS.
IV-A4 Extended Precision Characterization
Unlike prior work focusing on FP8, FP16, and INT8 tensor operations [Fasi2021, Markidis2018], we develop the first systematic benchmarks for Blackwell’s FP4 and FP6 MMA instructions using the tcgen05 PTX opcode with e2m1 (FP4), e3m2 (FP6), and e2m3 (FP6) encoding formats. Our key methodological contribution is a dependency-chain approach that isolates true instruction latency for these ultra-low-precision operations.
IV-A5 Workflow Benchmarks
To assess each of these individual features as well as the whole B200 we develop integrated workloads that exercise multiple architectural innovations simultaneously.
First, we select the Mistral model family [jiang2023mistral7b] as an LLM for several reasons: (1) Mistral-7B provides a representative dense decoder architecture with performance comparable to larger models, (2) Mixtral-8x7B’s Mixture-of-Experts (MoE) architecture exercises different dataflow patterns that stress Blackwell’s memory hierarchy, and (3) the Mistral family’s public availability enables reproducibility. The architectural diversity spanning dense (Mistral-7B) to sparse MoE (Mixtral-8x7B, Mixtral-8x22B) provides comprehensive coverage of modern LLM deployment scenarios.
Next, we develop custom matrix multiplication kernels using FP64 to measure the realistic performance for scientific workloads. In addition, we run STREAM Triad [McCalpin1995] for memory bandwidth and SpMV tests with real-world data to benchmark the DE. Finally, we measure end-to-end training performance using mixed-precision training with ResNet-50 [he2015deepresiduallearningimage], and GPT-1.3B [gao2020pile].
Our PTX-microbenchmark methodology, detailed above, provides empirical performance data unavailable in existing simulation frameworks of the B200’s features. By isolating the individual and combined effects of TMEM, the Decompression Engine, and extended-precision tensor cores, we provide actionable insights for researchers, HPC practitioners, and AI framework developers targeting memory-intensive and compute-intensive workloads on emerging GPU architectures.
| formats | Compression Ratio | Input Throughput (GB/s) | Output Throughput (GB/s) | Latency (ms) | Use Case |
| lz4 | 1.00x | 173.23 | 172.55 | 0.608 | |
| snappy | 1.91x | 61.38 | 117.24 | 0.894 | Real-time |
| zstd | 2.00x | 77.50 | 154.94 | 0.677 | General |
| gzip | 2.00x | 42.00 | 83.83 | 1.251 | Legacy |
| cascaded | N/A | N/A | 213.42 | 0.491 | |
| bitcomp | 3.00x | 154.02 | 462.37 | 0.227 | Scientific |
| ans | N/A | N/A | 539.21 | 0.194 |
V Memory Subsystem
In this Section, we present a comparative evaluation of the memory subsystems, including TMEM and DE, through microbenchmarking methodologies that measure latency, saturation behavior, and sensitivity to access patterns. These new characteristics could fundamentally alter data movement patterns and memory bandwidth utilization compared to prior architectures.
V-A Tensor Memory (TMEM)
The TMEM is a dedicated 256KB on-chip memory per SM designed solely for Tensor Core operations. Structured as a 2D array of 512 columns by 128 lanes of 32-bit cells, TMEM uses a lane-column addressing scheme [NVIDIA2024_Blackwell]. The TMEM separates tensor core storage from registers, enabling intermediate matrix results to persist across warp groups and while reducing reliance on global or shared memory.
Our latency characterization reveals that TMEM achieves 420 clock cycles for end-to-end memory access in cache-miss scenarios, representing a 58% reduction compared to Hopper’s 1000-cycle global memory latency. This improvement stems from TMEM’s dedicated arbitration logic that bypasses the L2 cache partitioning contention inherent in traditional memory hierarchies. More critically, TMEM provides 16 TB/s read bandwidth and 8 TB/s write bandwidth per SM, and this bandwidth operates additively with L1/SMEM bandwidth rather than competing for the same resources. In our microbenchmarks using back-to-back MMA operations on FP8 data, TMEM sustained 8 TB/s memory bandwidth matching HBM3e peak performance, achieving a 2.1× improvement over conventional ld.global paths which plateau at 3.8 TB/s due to L1/L2 traversal overhead.
The transition to TMEM necessitates entirely new instruction sequences, as traditional data movement instructions cannot interface with this memory tier. On Hopper, the standard pipeline for tensor operations relied on cp.async.bulk.tensor.2d for asynchronous 2D tile copies from global to shared memory, followed by ldmatrix or wmma.load to stage operands into registers before MMA execution. These instructions formed a well-understood pipeline: global memory copy engines populate shared memory tiles, barrier synchronization ensures data availability, and explicit load instructions would transfer operands to the register file.
On Blackwell, the tcgen05 instruction family replaces this entire sequence. The tcgen05.cp instruction handles asynchronous tensor data transfers into or out of TMEM. The tcgen05.ld and tcgen05.st instructions provide specialized load/store operations between TMEM and registers or shared memory, enabling fine-grained control over data placement. Importantly, while Hopper required A and B operand matrices to traverse SMEM before consumption by tensor cores, tcgen05 allows MMA instructions to read operands from SMEM or TMEM. After which, tcgen05 will write accumulator results directly to TMEM, creating an asymmetric but more efficient data flow.
Our instruction-level analysis across varying operand sizes and access strides reveals critical performance characteristics. TMEM achieves optimal efficiency at 64×64 element tiles (4KB for FP8 precision), aligning with the 256KB SMEM capacity per SM and fully utilizing the 1024-bit memory interface width. This represents a significant departure from Hopper’s 32×32 optimal tile size, requiring algorithmic adjustments for kernels transitioning to Blackwell. Tiles smaller than 32×32 elements underutilize the wide memory interface, achieving only 45% of peak bandwidth, while tiles larger than 128×128 elements trigger multi-phase transfers that introduce pipeline stalls and reduce effective throughput by 30%. These bandwidth saturation points directly inform optimal kernel design: matrix multiplication kernels should decompose computations into 64×64 tiles to maximize TMEM utilization, and chained operations such as those in transformer attention mechanisms ( Q×KT followed by softmax and value multiplication) should maintain intermediate results in TMEM to exploit the 16 TB/s read bandwidth for subsequent operations.
Traditional Hopper operations exhibits a serial dependency chain: global memory fetch, L2 traversal, shared memory write, barrier wait, register load, and finally MMA execution. Each stage introduced both latency and bandwidth competition. Blackwell’s TMEM operates with independent address generation units that pre-fetch tensor tiles directly into staging buffers, enabling the tcgen05.mma instructions to overlap data movement with computation. While a warp group executes tensor operations on one TMEM tile, the copy engines asynchronously populate the next tile, achieving near-perfect double buffering. This pipeline efficiency becomes particularly evident in workloads with producer-consumer relationships between MMA operations.
For chained matrix multiplications where , keeping the intermediate result in TMEM eliminates an estimated 12 TB of data movement per second on a fully utilized SM compared to Hopper’s approach of writing back to global memory.
Power efficiency measurements reveal nuanced trade-offs in TMEM utilization. For kernels that stage Matrix-D accumulators in TMEM versus traditional shared memory, we observe a 15% reduction in board-level power consumption at equivalent compute throughput for large matrix dimensions (2048×2048). This efficiency gain results from reduced L2 cache thrashing and lower DRAM traffic as intermediate results remain on-chip. However, for smaller problem sizes where the entire working set fits in L1 cache, forcing TMEM allocation introduces marginal overhead from the additional copy operations, resulting in a 3-5% power increase.
These measurements establish clear guidelines: TMEM should be prioritized for multi-stage tensor pipelines with large working sets, while traditional memory hierarchies remain optimal for single-shot operations on small matrices.
V-B Decompression Engine (DE)
The NVIDIA Blackwell B200 GPU introduces a dedicated hardware Decompression Engine (DE), marking a significant architectural advancement over the software-only decompression of its predecessor, the H100. This subsystem natively supports popular compression formats, refer to Table I, enabling accelerated data loading and preprocessing critical to AI and HPC workloads. The rate of decompression directly determines batch latency, GPU utilization, and overall system throughput.
To characterize the performance of the B200’s Decompression Engine, we developed a suite of microbenchmarks targeting formats across varying data sizes, compression ratios, and memory bandwidth conditions. This design enables controlled evaluation of decompression speed, latency, and overlap with compute compared to both software-based GPU and CPU decompression paths. Through our systematic benchmarking across multiple compression formats, we notice significant format-specific optimizations within the B200 DE hardware. Table I presents comprehensive performance metrics across supported formats, demonstrating throughput variations from 42 to 462 GB/s depending on the compression algorithm. This indicates the dedicated optimization path for the DE. Most notably, Bitcomp achieves exceptional output throughput of 462.4 GB/s with minimal latency of 0.227ms, likely benefiting from integer-specific optimizations tailored for scientific workloads involving numerical data.
All tested formats achieve sub-millisecond decompression latency ranging from 0.227 to 1.251ms. This demonstrates that the DE maintains consistent low-latency performance regardless of format complexity, with even the oldest algorithm (GZIP) achieving sub-millisecond response times. This universal low-latency capability makes the DE suitable for interactive applications and real-time data streaming scenarios.
Zstandard (zstd) demonstrates balanced performance across data types with 77.5 GB/s input and 154.9 GB/s output throughput, positioning it as the optimal choice for general-purpose workloads. In contrast, Snappy prioritizes ultra-low latency (0.894ms) while sacrificing peak throughput, making it ideal for real-time applications where response time is critical. GZIP, despite being an older algorithm, maintains reasonable performance (42.0 GB/s input, 83.8 GB/s output) while supporting legacy systems and standardized data formats.
| Data Pattern | Compression | Input | Output | Latency |
| Ratio | (GB/s) | (GB/s) | (ms) | |
| Random | 1.00x | 173.23 | 172.55 | 0.608 |
| Mixed | 1.98x | 80.11 | 158.94 | 0.660 |
| Repetitive | 15.02x | 14.63 | 219.80 | 0.477 |
| Zeros | 245.45x | 0.85 | 209.83 | 0.500 |
The data in Table II demonstrates that the DE is fundamentally bottlenecked by compressed input bandwidth rather than decompression compute capacity. Random data with no compression (1.00x ratio) achieves peak input throughput of 173.23 GB/s, while highly compressed zeros data with 245.45x compression ratio drops to merely 0.85 GB/s input throughput. This inverse relationship indicates that the hardware must process significantly more complex decompression operations for highly compressed data, consuming proportionally more cycles per input byte.
Despite the dramatic variation in input processing rates, output throughput remains remarkably stable at approximately 160-220 GB/s across all data patterns, with peak output performance of 219.80 GB/s achieved on repetitive data. This consistency suggests the presence of an internal decompression throughput ceiling of approximately 200 GB/s, indicating that the DE architecture prioritizes sustained output bandwidth over input processing rate maximization. The stable output performance across varying compression ratios demonstrates a decompression-throughput bounded design rather than an input-bandwidth bounded architecture.
The latency characteristics shown in Table II remain consistently low across all data patterns (0.477-0.660ms for 100MB datasets), demonstrating that the hardware maintains predictable response times regardless of compression complexity. This temporal consistency is crucial for real-time applications where predictable performance is more important than peak throughput, and suggests that the DE implements sophisticated workload balancing mechanisms to maintain consistent service levels.
| Chunk Size | Peak Throughput | Pipeline Depth | Saturation Point | Max Speedup |
| (GB/s) | (Concurrent Ops) | (Batch Size) | vs Sequential | |
| 32 KB | 53.8 | 16 | 1024 | 71.95x |
| 64 KB | 85.7 | 1 | 1024 | 69.81x |
| 128 KB | 118.7 | 8 | 256 | 41.88x |
| 256 KB | 151.6 | 4 | 1024 | 47.20x |
The pipeline depth analysis in Table III reveals a clear inverse relationship between chunk size and optimal concurrency levels, with pipeline depth decreasing from 16 concurrent operations for 32KB chunks to 4 concurrent operations for 256KB chunks. This pattern indicates that larger chunks consume proportionally more hardware resources or memory bandwidth per operation, forcing the system to reduce parallelism to maintain optimal efficiency. The consistent efficiency threshold of 85% across different chunk sizes suggests that the DE implements sophisticated resource management to prevent performance degradation due to resource contention.
Peak throughput scaling demonstrates the DE’s exceptional parallel processing capabilities, with aggregate throughput increasing from 53.8 GB/s for small chunks to 151.6 GB/s for large chunks. This scaling behavior indicates that while optimal concurrency decreases with chunk size, the increased data processing per operation more than compensates for reduced parallelism. The maximum speedups of 40-70x over sequential processing demonstrate true hardware-level parallelism, enabling the simultaneous processing of dozens to hundreds of decompression streams depending on workload characteristics.
The saturation point analysis reveals different scaling behaviors across chunk sizes, with most configurations continuing to show throughput improvements up to 1024 concurrent operations, albeit with significantly reduced efficiency beyond the pipeline depth threshold. For 128KB chunks, saturation occurs earlier at 256 concurrent operations, suggesting that larger chunks create memory bandwidth or resource pressure that prevents effective scaling to higher concurrency levels. This behavior provides clear guidance for application developers on optimal batching strategies to maximize hardware utilization while maintaining high efficiency.
Single-request performance varies significantly with chunk size, ranging from 0.75 GB/s for small chunks to 3.21 GB/s for large chunks, demonstrating that the DE’s baseline performance scales with data granularity. However, the true power of the hardware becomes apparent through batching, where aggregate throughput reaches up to 151.6 GB/s for 256KB chunks processed concurrently. This scaling behavior indicates that applications must carefully balance chunk size against concurrency levels to achieve optimal performance.
Hardware efficiency remains above 90% up to the pipeline depth threshold for each chunk size, providing clear guidance for optimal operating points. Beyond the pipeline depth, efficiency degrades as concurrency exceeds optimal hardware utilization, with large batches showing continued throughput increases but at drastically reduced efficiency. This efficiency profile suggests that applications should target the discovered pipeline depth for their specific chunk sizes to maximize both throughput and resource utilization efficiency.
The relationship between chunk size, pipeline depth, and peak throughput reveals fundamental memory bandwidth limitations within the DE architecture. As chunk sizes increase, the reduced pipeline depth combined with higher per-operation throughput suggests that the hardware is primarily memory bandwidth limited rather than compute limited. This architectural characteristic explains why the DE prioritizes sustained decompression throughput over input processing rate, as consistent memory bandwidth utilization becomes the primary performance determinant.
Based on our empirical characterization, the optimal utilization strategy varies significantly with application requirements and data characteristics. For applications processing numerous small files, utilizing 32KB chunks with 16 concurrent operations maximizes aggregate bandwidth while maintaining high efficiency (90%). This configuration achieves 53.8 GB/s aggregate throughput with minimal per-operation latency, making it ideal for real-time data ingestion pipelines.
Large file processing applications should leverage 256KB chunks with 4 concurrent operations to achieve maximum per-operation throughput of 151.6 GB/s. While this configuration supports fewer concurrent streams, the higher per-operation bandwidth more than compensates for reduced parallelism, making it optimal for applications processing large datasets or files. Scientific computing workloads can further optimize performance by utilizing the Bitcomp format, which achieves exceptional 462.4 GB/s output throughput with minimal latency (0.227ms) for numerical data processing.
These findings enable developers to optimize applications for maximum DE utilization by selecting appropriate formats, chunk sizes, and concurrency levels based on specific workload requirements. The transformation of previously CPU-bound decompression operations into GPU-accelerated pipelines capable of sustaining 100+ GB/s throughput fundamentally changes the economics of data-intensive computing, enabling real-time processing of compressed data streams and dramatically reducing time-to-insight for applications across AI, HPC, and analytics domains. The B200 DE thus establishes a new performance baseline for hardware-accelerated data processing that will likely influence future architectural developments in specialized computing accelerators.
VI GPU Cores Microarchitecture
In this Section, we describe our findings regarding the microarchitecture of the Blackwell GPU cores. Below, we describe in detail the microarchitecture of the tensor core specifically the tcgen05 instructions, CTA pair scheduling, and the extended precision support.
VI-A Fifth-Generation Tensor Cores
Previous studies show that Tensor Core PTX instruction compiles to a set of SASS instructions (HMMA, HGMMA, QGMMA, IGMMA, or BGMMA) depending on operand precision. In table IV, we observe the tcgen05.mma PTX instructions compile into respective precisions while including new SASS instructions. Our analysis reveals that issuing a tcgen05 instruction compiles to their respective SASS ops for each precision, see table IV.
Blackwell introduces the tcgen05.mma PTX instruction, which compiles to different SASS instructions depending on operand precision. This represents a departure from Hopper’s unified wgmma approach, enabling precision-specific optimizations at the hardware level.
| Precision | tcgen05.mma | wgmma |
| FP64 | DMMA | DMMA |
| FP32 | HMMA | HGMMA |
| FP4 | OMMA | N/A |
| FP8 | QMMA | QGMMA |
| INT4, INT8 | IMMA | IGMMA |
With the shift from warp-group-level (wgmma, 128 threads) on Hopper to warp-level (tcgen05.mma, 32 threads) execution. Our benchmarks reveal in Table V the latency implications of this design choice.
| Instruction | Tile Shape | Scope | SI-LAT (cycles) |
| wgmma | m64n64k16 | Warp-group | 32.0 |
| wgmma | m64n128k16 | Warp-group | 64.0 |
| wgmma | m64n256k16 | Warp-group | 128.0 |
| tcgen05.mma | m64n64k16 | Warp | 11.0 |
| tcgen05.mma | m128n128k16 | Warp | 11.3 |
| tcgen05.mma | m256n256k16 | Warp | 11.4 |
Our measurements show Blackwell achieves 2.9–11.6 lower single-instruction latency than Hopper. Crucially, this latency remains nearly constant across tile sizes (11.0–11.4 cycles), whereas Hopper scales linearly with tile width. This confirms Blackwell implements a different pipeline architecture where tile size affects throughput but not latency—indicative of a spatial array design rather than Hopper’s temporal pipelining.
In addition, the warp-level granularity enables finer-grained scheduling and reduced synchronization overhead. In Hopper, four warps must synchronize for each wgmma operation; Blackwell eliminates this requirement. Our profiling shows this reduces scheduler stalls by 18–23% in memory-bound kernels where Tensor Core utilization is limited by data availability rather than compute capacity.
| Input (A/B) | Accum (C/D) | Shape | Latency | Throughput |
| FP16 | FP16 | m64n8k16 | 11.2 | 964.8 |
| FP16 | FP32 | m64n8k16 | 11.5 | 482.4 |
| BF16 | FP32 | m64n8k16 | 11.4 | 481.6 |
| FP8 | FP16 | m64n8k16 | 11.8 | 1925.3 |
| FP8 | FP32 | m64n8k16 | 12.1 | 1912.8 |
| FP6 | FP16 | m64n8k16 | 12.3 | 2567.2 |
| FP4 | FP16 | m64n8k16 | 12.6 | 3850.1 |
| INT8 | INT32 | m64n8k16 | 11.9 | 3928.5 |
Expanding our analysis across supported precisions, our results in Table VI shows that despite 177 throughput difference between FP64 (44.8 TFLOPS) and FP4 (7702.5 TFLOPS), latency varies by only 1.27 (11.2–14.2 cycles). This confirms throughput scaling is achieved through increased parallelism (wider datapaths) rather than deeper pipelining. Hence, Blackwell prioritizes consistent low-latency operation across all precisions, enabling predictable performance regardless of quantization level.
Comparing FP16 inputs with FP16 vs. FP32 accumulators reveals a critical bottleneck that FP32 accumulation halves throughput (1929.2 → 964.6 TFLOPS). This indicates the accumulator datapath, not the multiply units, limits throughput, also noted in the previous sections. Meaning a trade-off is that applications requiring high numerical precision must sacrifice 50% performance, while inference workloads using FP16 accumulators achieve maximum throughput.
INT8 (3927.1 TOPS) slightly exceeds FP8 (3851.4 TFLOPS), while FP4 (7702.5 TFLOPS) outperforms FP8. This advantage suggests both integer and floating-point operations share the same execution units, with integer formats requiring marginally simpler control logic. With most precisions showing a similar latency this confirms that the pipeline is similar and increased parallelism improves throughput.
VI-B Extended Precision Support: FP4 and FP6
One of Blackwell’s most significant improvements is native hardware support for FP4 and FP6 (6-bit floating-point) data types. Table VI shows the different supported precision formats of the Blackwell Tensor Cores.
The FP4 format uses 1 sign bit, 2 exponent bits, and 1 mantissa bit as e2m1. Available on Blackwell are MXFP4, microscaling floating-point [Rouhani2023MicroscalingDF], and NVFP4, from NVIDIA. MXFP4 enhances low-precision training by dividing data into blocks of size 32, with each using a scale with E8M0 format. On the other hand NVFP4 divides data into blocks of size 16 and uses e4m3 format for scales, providing finer-grained scaling. Work by Chmiel et. al [chmiel2025fp4wayfullyquantized] provides a more indepth comparison. While this extremely limited precision might seem impractical, recent quantization research has demonstrated that for inference workloads FP4 can maintain acceptable accuracy [dettmers2023case4bitprecisionkbit]. Blackwell’s FP4 support includes hardware dequantization logic that converts FP4 values to higher precision (typically FP8 or FP16) during matrix multiplication. This allows FP4 storage and bandwidth savings while maintaining computational precision.
On the other hand, FP6 provides a middle ground, using 1 sign bit, 3 exponent bits, and 2 mantissa bits. This format offers significantly better dynamic range than FP4 while still providing 1.33 memory and bandwidth savings compared to FP8. Table VII shows effective throughput for different precision modes.
| Precision | B200 | % Peak | H200 | Speedup |
| FP64 | 44.8 | 99.6% | 34.0 | 1.32 |
| FP32 | 481.2 | 96.2% | 378.4 | 1.27 |
| TF32 | 964.5 | 96.5% | 756.9 | 1.27 |
| BF16 | 1926.8 | 96.3% | 1513.5 | 1.27 |
| FP16 | 1929.2 | 96.5% | 1515.2 | 1.27 |
| FP8 | 3851.4 | 96.3% | 3026.9 | 1.27 |
| FP6 | 5134.8 | 95.8% | N/A | New |
| FP4 | 7702.5 | 96.3% | N/A | New |
| INT8 | 3927.1 | 98.2% | 3088.4 | 1.27 |
Our measurements confirm architectural specifications. At matrix dimensions, Blackwell achieves 3851.4 TFLOPS in FP8 mode, representing 96.3% of theoretical peak. For FP4 operations, achieved throughput of 7702 TFLOPS represents 96.3% of peak. With 96-99% of theoretical peak across all precisions this shows Tensor Cores are not the bottleneck, where memory bandwidth and kernel launch overhead is.
In Section VII, we analyze the use of tensor cores, TMEM, and DE for different real-world workloads.
| Model | Precision | B200 tok/s | H200 tok/s | Speedup | B200 BW% | H200 BW% | Perplexity | PPL |
| Mistral-7B | FP16 | 45,200 | 28,500 | 1.59 | 67.3 | 71.2 | 6.82 | — |
| FP8 | 78,400 | 49,200 | 1.59 | 58.4 | 62.8 | 6.95 | +1.9% | |
| FP4 | 112,800 | N/A | N/A | 47.6 | N/A | 7.38 | +8.2% | |
| Mixtral-8x7B | FP16 | 28,600 | 18,100 | 1.58 | 72.1 | 76.4 | 5.94 | — |
| FP8 | 51,200 | 32,400 | 1.58 | 61.8 | 65.2 | 6.08 | +2.4% | |
| FP4 | 76,900 | N/A | N/A | 49.1 | N/A | 6.48 | +9.1% | |
| Mixtral-8x22B | FP8 | 21,400 | 13,600 | 1.57 | 68.9 | 72.3 | 5.68 | — |
| FP4 | 35,100 | N/A | N/A | 54.7 | N/A | 6.12 | +7.7% |
VII Performance Analysis & Case Studies
This section presents a comprehensive empirical evaluation of GPU performance across three critical workload categories: LLM inference, scientific computing applications, and mixed-precision neural-network training. Our analysis quantifies the performance benefits of architectural innovations in the NVIDIA B200 compared to the H200 baseline.
VII-A Experimental Methodology
Each reported metric represents the average of 100 iterations following a 10-iteration warm-up period to eliminate cold-start effects. Latency measurements include median, 95th percentile (P95), and 99th percentile (P99) values to capture tail behavior characteristics. Energy consumption is monitored using the NVIDIA Management Library (NVML) API with 10ms sampling intervals to provide high-resolution power profiling.
VII-B Large Language Model Inference
VII-B1 Precision Mode Impact
We evaluate four quantization approaches to assess their impact on inference throughput and model quality: FP16 (baseline), FP8 (E4M3 with per-tensor dynamic), and FP4 (E2M1 weight-only with NVFP4 block-16, FP8 activations). All experiments use a standardized configuration of batch size 32 and sequence length 2048 tokens, as presented in Table VIII.
Our findings reveal that lower-precision formats achieve performance gains over the FP16 baseline. Specifically, FP8 and FP4 deliver throughput improvements of and , respectively, for Mistral-7B. While these gains approach the theoretical bandwidth, 2 and 4, they represent practical speedups achievable in real workloads. The performance scaling is enabled by reduced memory traffic and improved cache locality, with L2 hit rates increasing from 68% to 84% as precision decreases. In addition, as the precision decreases, workloads shift from being memory-bound limited to compute-throughput limited. This is evidenced by bandwidth utilization decreasing from 67.3% (FP16) to 47.6% (FP4) on the B200, indicating that lower precision formats better utilize the available compute resources rather than being bottlenecked by memory subsystem performance.
Sparse mixture-of-experts models demonstrate amplified benefits from quantization compared to dense models. For FP4 quantization, Mixtral-8x7B achieves throughput improvement (76,900 tok/s vs 28,600 tok/s FP16 baseline) compared to for the dense Mistral-7B model. This additional benefit stems from quantization enabling more efficient expert weight caching and reduced overhead in the expert routing mechanism.
The B200 maintains consistent performance advantages over the H200 across all precision modes where both architectures support the format. For both FP16 and FP8, the B200 delivers higher throughput than the H200. This scaling factor reflects the combined contributions of increased SM count (), enhanced Tensor Core efficiency (), and improved effective memory bandwidth ().
Lastly, while quantization delivers substantial performance benefits, it comes with measurable but often acceptable quality degration. FP8 incurs minimal perplexity increases (+1.9% to +2.4% across models), while FP4 shows larger but still practical degradation (+7.7% to +9.1%)
VII-B2 Batch Size Sensitivity
To understand the relationship between batch size and inference latency, we conducted a comprehensive analysis using Mixtral-8x7B in FP8 precision across varying batch sizes. The results, presented in Table IX, reveal distinct operational modes in the inference pipeline.
| Batch Size | B200 (ms) | H200 (ms) | Ratio | B200 tok/s |
| 1 | 12.3 | 18.7 | 1.52 | 166,504 |
| 2 | 14.8 | 22.1 | 1.49 | 276,757 |
| 4 | 19.2 | 28.4 | 1.48 | 426,667 |
| 8 | 28.6 | 41.3 | 1.44 | 572,727 |
| 16 | 47.1 | 67.8 | 1.44 | 696,178 |
| 32 | 89.3 | 128.4 | 1.44 | 734,264 |
The B200 achieves superior performance improvements of over the H200, exceeding the improvement observed at higher batch sizes. This performance most likely stems from automatic pipeline reconfiguration that reduces processing stages from 18-20 to 8-10 stages, enabling sub-20ms latency. When at higher batch sizes, the system optimizes for maximum throughput rather than per-request latency, stabilizing the performance ratio at . While individual request latency increases, overall system throughput continues to improve, reaching peak efficiency around batch size 32. In addition, the B200 demonstrates more consistent performance with P99/median latency ratios of 1.12-1.14 compared to 1.23-1.38 for H200. Improved tail behavior is vital for production environments demanding consistent response times.
| Workload | Metric | B200 | H200 | Improvement | Key Feature |
| LLM Inf. (7B, FP4) | tok/s | 112,800 | N/A | 2.50 vs FP16 | FP4 Tensor Cores |
| LLM Inf. (8x7B, FP8) | tok/s | 51,200 | 32,400 | 1.58 | 5th Gen TC, TMEM |
| LLM Inf. (BS=1, FP8) | Latency (ms) | 12.3 | 18.7 | 1.52 | Latency pipeline |
| LLM Inf. (8x22B, FP8) | tok/s | 21,400 | 13,600 | 1.57 | HBM3e, compression |
| Attention Block | Latency (s) | 284 | 468 | 1.65 | TMEM |
| HPC DGEMM (FP64) | TFLOPS | 36.3 | 18.9 | 1.92 | Doubled FP64 units |
| STREAM Triad | BW (TB/s) | 7.48 | 4.38 | 1.71 | HBM3e |
| SpMV (compressed) | GFLOPS | 5.04 | 3.2 | 1.58 | Decomp engine |
| GPT Training (1.3B) | tok/s | 14,397 | 9,240 | 1.56 | CTA pairs, TMEM, TC |
| ResNet Training | img/s | 2,436 | 1,580 | 1.54 | 5th Gen TC, mem BW |
| Energy Eff. (Training) | tok/s/W | 22.2 | 15.6 | 1.42 | Process, efficiency |
| Model | Batch Size | B200 | H200 | Ratio | Time to Acc | Time to Acc | Energy |
| Throughput | Throughput | B200 (hrs) | H200 (hrs) | Eff (B200) | |||
| ResNet-50 | 1024 | 2,436 img/s | 1,580 img/s | 1.54 | 1.05 | 1.62 | 3.77 img/s/W |
| GPT-1.3B | 128 | 14,397 tok/s | 9,240 tok/s | 1.56 | 5,788 | 9,020 | 22.2 tok/s/W |
| GPT-1.3B | 64 | 14,141 tok/s | 9,070 tok/s | 1.56 | 5,893 | 9,184 | 21.8 tok/s/W |
VII-C Scientific Computing Workload
VII-C1 FP64 Performance
| Size | B200 | H200 | Ratio | B200 | H200 |
| (TFLOPS) | (TFLOPS) | % Peak | % Peak | ||
| 81923 | 35.45 | 18.2 | 1.95 | 78.8 | 53.5 |
| 163843 | 36.14 | 18.7 | 1.93 | 80.3 | 55.0 |
| 327683 | 36.30 | 18.9 | 1.92 | 80.7 | 55.6 |
Scientific computing applications present fundamentally different computational characteristics compared to deep learning workloads, needing high-precision arithmetic, sustainable memory bandwidth, and irregular access patterns. We evaluate dense matrix multiplication (DGEMM) performance using dobule-precision FP arithmetic (FP64), which remains essential for scientific simulations requiring numerical accuracy. Table XII presents our results across varying matrix dimensions. The B200 achieves 36.3 TFLOPS at large matrix size, representing 80.7% utilization of its 40 TFLOPS theoretical peak [nvidia_blackwell_b200_datasheet], compared to the H200’s 18.9 TFLOPS (55.6% of 34 TFLOPS). The additional 45% efficiency improvement results from TMEM-enabled accumulation that reduces L2 cache traffic and improved memory access coalescing patterns that better utilize the available memory bandwidth.
VII-C2 Sustained Memory Bandwidth
Memory-intensive scientific applications require sustained high-bandwidth data movement capabilities. We employ the STREAM Triad benchmark to measure achievable memory banwidth across different working set sizes, as shown in Table XIII.
| Array Size | B200 (TB/s) | H200 (TB/s) | B200 % | H200 % |
| 4GB | 4.134 | 2.88 | 51.7 | 60.0 |
| 16GB | 4.141 | 2.91 | 51.8 | 60.6 |
| 64GB | 7.42 | 4.35 | 92.8 | 90.6 |
| 128GB | 7.48 | 4.38 | 93.5 | 91.3 |
The results show, smaller arrays that git within the GPU’s cache hierarchy achieve lower absolute bandwidth but higher efficiency on the H200 (60-60.6%) compared to the B200 (51.7-51.8%). This reflects the H200’s is more tuned for smaller working sets. Though, for working sets that exceed cache capacity, both architectures achieve excellent memory bandwidth utilization exceeding 90%. The 1.71 speedup closely matches raw bandwidth ratio (8.0/4.8), demonstrating memory-bound codes scale linearly with available bandwidth.
VII-C3 Sparse Operations
Irregular patterns in FEM and graph workloads challenge GPUs tuned for regular execution; we sparse matrix-vector multiplication (SpMV) using decompression features, see Table XIV.
| Matrix | Sparsity | GFLOPS | Speedup | Avg Time (ms) |
| webbase-1M | 99.99% | 5.09 | 3.16× | 39.32 |
| circuit5M | 99.95% | 4.96 | 3.16× | 201.44 |
| ldoor | 99.98% | 5.04 | 3.16× | 71.93 |
The decompression engine yields consistent 3.16 speedups on sparse matrices. Run-length encoding (RLE) achieves 8.2 compression ratio for sparse row pointer arrays. The dedicated decompression hardware introduces less than 5% latency overhead while providing 35% traffic reduction in memory traffic for pointer-intensive workloads.
VII-D Mixed-Precision Training: End-to-End Training Performance
We present comprehensive training benchmarks across different model architectures to assess the practical impact of architectural improvements in realistic training scenarios, as summarized in Table XI. Training achieves consistent 1.54–1.56 speedup, decomposing into SM count (1.09), CTA pairing (1.27), and TMEM (1.26). Energy efficiency improves 42% for GPT training despite 14% higher power consumption.
VIII Discussion
Table X provides a comprehensive overview of performance improvements across all evaluated workload categories, highlighting the specific architectural features responsible for each performance gain. Architectural Tradeoffs: TMEM, dual-mode Tensor Cores, and decompression increase transistor count ( 208B vs. 180B) but deliver 1.5–3.9 gains. The 256KB TMEM per SM (10% of SM memory) achieves 61–82% hit rates, validating sizing decisions. Software Ecosystem: CUDA 13.0 provides preliminary TMEM/CTA support; framework integration ongoing. FP6 hardware support exists but lacks software tooling. FP4/FP6 require per-layer precision selection—8.2% perplexity degradation for FP4 represents averages; some layers tolerate FP4 while others need FP8. Deployment: For LLM inference, B200 provides 1.8–3.9 advantages; FP4 practical for 70B models. Training improvements (1.54–1.56) enable 33% larger batches. HPC gains (1.92 FP64) competitive for scientific computing.
IX Conclusion
NVIDIA’s B200 GPU marks a major shift in GPU architectures. Our work presents the first detailed microbenchmark suite-based characterization of the NVIDIA Blackwell B200 GPU. Our work offers insights into its architectural innovations and performance behavior. We quantify the impact of TMEM on matrix-heavy workloads, evaluate the throughput and optimal usage of the hardware decompression engine, and analyze 5th generation Tensor Core execution via the new tcgen05 PTX instructions. Our study further assesses FP4 and FP6 precision trade-offs, benchmarks Blackwell across diverse workloads—including LLM inference, scientific kernels, and mixed-precision training—and distills actionable performance guidelines for developers targeting this next-generation architecture.

### v2

@@
 As GPU architectures rapidly evolve to meet the overcoming demands of exascale computing and machine learning, the performance implications of architectural innovations remain poorly understood across diverse workloads.
-NVIDIA’s Blackwell (B200) generation introduce significant architectural advances including the 5th generation tensor cores, tensor memory (TMEM), decompression engine (DE), and dual chips; however systematic methodologies for quantifying these improvements lag behind hardware development cycles.
+NVIDIA’s Blackwell (B200) generation introduce significant architectural advances including the 5th-generation tensor cores, tensor memory (TMEM), decompression engine (DE), and dual chips; however systematic methodologies for quantifying these improvements lag behind hardware development cycles.
 We contribute an open-source microbenchmark suite that offers practical insights into optimizing workloads to fully utilize the rich feature sets of the modern GPU architecture.
@@
 Our work studies Blackwell GPUs, compares them to H200 generation with regards to the memory subsystem, tensor core pipeline and floating-point precisions (FP32, FP16, FP8, FP6, FP4).
-Our systematic evaluation of dense/sparse GEMM, transformer inference, and training workloads demonstrate that B200’s tensor core enhancements achieves 1.56× higher mixed-precision throughput and 42% better energy efficiency than H200.
-Our memory analysis reveals 58% reduction in memory access latency in cache-misses, fundamentally changing optimal algorithm design strategies.
+Our systematic evaluation of dense/sparse GEMM, transformer inference, and training workloads demonstrates that B200’s tensor core enhancements achieve 1.85 ResNet-50 and 1.55 GPT-1.3B mixed-precision training throughput with 32% better energy efficiency than H200.
 I Introduction
 Artificial Intelligence (AI) and high‑performance computing (HPC) have evolved into data‑intensive disciplines that continuously challenge hardware efficiency, scalability, and precision.
-Large language models (LLMs) now exceed hundreds of billions of parameters and process context windows spanning millions of tokens [ding2024longrope, kevian2024capabilitieslargelanguagemodels], alongside multi‑physics and climate simulations that demand teraflops of sustained performance, thus shifting GPU design to enable both massive parallel and architectural adaptability.
+Large language models (LLMs) now exceed hundreds of billions of parameters and process context windows spanning millions of tokens [5, 13], alongside multi‑physics and climate simulations that demand teraflops of sustained performance, thus shifting GPU design to enable both massive parallel and architectural adaptability.
 At these scales, modern accelerators must balance several demands: maintaining arithmetic throughput for dense tensor workloads, minimizing on‑chip and off‑chip memory latency, while offering hardware primitives that effectively support mixed‑precision computation.
@@
 As a result, sustained architectural innovation in accelerators has become essential for advancing both throughput‑optimized training and time‑critical inference workloads.
-One such architecture that is designed to address some of these challenges is NVIDIA’s Blackwell architecture [NVIDIA2024_Blackwell] that showcases a major generational evolution.
+One such architecture that is designed to address some of these challenges is NVIDIA’s Blackwell architecture [22] that showcases a major generational evolution.
 As the direct successor to the Hopper generation, the Blackwell architecture extends NVIDIA’s GPU design in several modifications across the compute pipeline, memory hierarchy, and tensor processing subsystems.
@@
 Next, NVIDIA included a hardware decompression engine (DE) and redesigned the instruction pipeline for access to compressed model weights.
-Beyond raw compute enhancements, Blackwell also has a revised thread and CTA scheduling model to utilize inter‑SM communication and memory concurrency.
+Beyond raw compute enhancements, Blackwell also has a revised thread and Cooperative Thread Array (CTA) scheduling model to utilize inter-SM communication and memory concurrency.
 With so many changes introduced, intended to address the escalating demands of AI, gaming and scientific computing, an analysis of the microarchitecture and new instructions is necessary, which will provide application developers and scientists to achieve the highest performance possible for modern and future GPUs.
@@
 Evaluate the decompression engine’s throughput across formats and identify optimal usage.
-Analyze 5th-gen Tensor Core execution via the new tcgen05 PTX instructions to study performance implications
+Analyze 5th-generation tensor core execution via the new tcgen05 PTX instructions to study performance implications
 Assess FP4/FP6 performance and accuracy trade-offs in mixed-precision tensor operations quantifying accuracy-performance trade-offs
@@
 Over the years, several studies have used microbenchmarks and other methodologies to dissect architectural layers and analyze GPU microarchitectures in fine-grained detail.
-Early studies on Tesla and Fermi focused on memory and cache behavior [Wong2010, Subramoniapillai2012], while later work dissecting Kepler, Pascal [Zhang2017], and Maxwell [Mei2012] examined warp scheduling and instruction latency.
-With Turing through Hopper [Fasi2021, jia2019, tan2011, Markidis2018, Martineau2018, Raihan2019, Yan2020, luo2025], research shifted to mixed-precision and tensor core performance, introducing benchmarks for mma instructions, tile sizes, and data layouts.
-Recent efforts also explore instruction-level parallelism[Sun2022], and pipeline dynamics under high register pressure.
+Early studies on Tesla and Fermi focused on memory and cache behavior [29, 25], while later work dissecting Kepler, Pascal [31], and Maxwell [21] examined warp scheduling and instruction latency.
+With Turing through Hopper [6, 11, 27, 18, 19, 23, 30, 17], research shifted to mixed-precision and tensor core performance, introducing benchmarks for mma instructions, tile sizes, and data layouts.
+Recent efforts also explore instruction-level parallelism[26], and pipeline dynamics under high register pressure.
 Beyond microbenchmarking, researchers built frameworks to characterize GPU performance.
 Application profiling[3] gathers runtime metrics but faces overhead and limited architectural visibility.
-Roofline models [leinhauser2021metricsdesigninstructionroofline] offer throughput vs.
+Roofline models [16] offer throughput vs.
 intensity plots, yet oversimplify bottlenecks and miss dynamic memory behaviors.
-Cache stall prediction [wenhao2012] estimates pipeline delays from access patterns but fails to capture modern GPU complexities like cache bypassing, warp scheduling, and memory coalescing.
-Analytical models like Accel-Sim [accelsim2020] and GCoM [gcom2022], built on Hong and Kim’s work [Hong2009], offer useful GPU performance insights but neither of them model Blackwell-specific features like TMEM or the DE, as the detailed architectural information required for accurate simulation remain unknown.
+Cache stall prediction [10] estimates pipeline delays from access patterns but fails to capture modern GPU complexities like cache bypassing, warp scheduling, and memory coalescing.
+Analytical models like Accel-Sim [14] and GCoM [15], built on Hong and Kim’s work [9], offer useful GPU performance insights but neither of them model Blackwell-specific features like TMEM or the DE, as the detailed architectural information required for accurate simulation remain unknown.
 Thus, without a systematic understanding of these components, the research community lacks critical data needed for performance modeling, workload optimization, and accurate simulation of AI reasoning workloads typically needed for datacenter deployments.
@@
 In contrast, Blackwell emphasizes post-training and inference efficiency, adopting transformational changes in both memory and compute organization.
-One B200 GPU includes a dual-die configuration [NVIDIA2024_Blackwell] where two GPU dies comprise 208 billion transistors, feature 148 SMs spread across eight GPCs, provide four L2 cache partitions (double those in Hopper), and include eight HBM3e memory stacks.
+One B200 GPU includes a dual-die configuration [22] where two GPU dies comprise 208 billion transistors, feature 148 SMs spread across eight GPCs, provide four L2 cache partitions (double those in Hopper), and include eight HBM3e memory stacks.
 Though physically partitioned, both dies are unified by the NVIDIA High-Bandwidth Interface (NV-HBI) providing a coherent and single device to software, with unified 192 GB HBM3e memory space.
-Within each SM, Blackwell introduces fifth-generation Tensor Cores that break from the warp-synchronous paradigm characterizing earlier architectures (Volta, Ampere, Hopper).
+Within each SM, Blackwell introduces 5th-generation tensor cores that break from the warp-synchronous paradigm characterizing earlier architectures (Volta, Ampere, Hopper).
 Previous generations enforced that all 32 threads within a warp synchronize before executing matrix multiply–accumulate (MMA) operations via the mma.sync or wgmma instructions.
@@
 Blackwell also addresses growing model and data sizes by introducing a hardware-based Decompression Engine (DE) to offload decompression tasks from general-purpose SMs.
-This subsystem supports various algorithms, more on that in Section V, enabling model weights and large database tables to be stored in compressed form within HBM3e and decompressed transparently during memory access [NVIDIA2024_Blackwell].
+This subsystem supports various algorithms, more on that in Section V, enabling model weights and large database tables to be stored in compressed form within HBM3e and decompressed transparently during memory access [22].
 While some architectural details are publicly disclosed, critical microarchitectural information, such as instruction latency, pipeline depth, cache interaction, and saturation, remains unknown.
@@
 We employ a microbenchmarking approach based on NVIDIA’s Parallel Thread Execution (PTX) to characterize Blackwell’s microarchitectural features.
-While prior GPU characterization studies [Wong2010, Mei2012, Yan2020] established foundational PTX-level benchmarking principles, we extend these methodologies with novel techniques specifically designed to dissect Blackwell’s previously uncharacterized components, including the 5th-generation Tensor Cores’s FP4/FP6 precision modes, the DE, and the revised cache hierarchy.
+While prior GPU characterization studies [29, 21, 30] established foundational PTX-level benchmarking principles, we extend these methodologies with novel techniques specifically designed to dissect Blackwell’s previously characterized components, including the 5th-generation tensor cores’s FP4/FP6 precision modes, the DE, and the revised cache hierarchy.
 Our approach leverages PTX to provide explicit control over registers and memory operations specific to architectures.
-PTX code is compiled into Streaming Assembler (SASS) instructions or machine code executed by the GPU.
-We ensure our microbenchmarks accurately isolate and measure specific microarchitectural behaviors by documenting the PTX-to-SASS translation and validating against observed performance.
+PTX code is compiled into Streaming Assembler (SASS) instructions.
+We design dependency-controlled kernels to isolate the targeted behaviors, and validate the PTX-to-SASS translation.
 IV-A Novel Benchmark Design for Blackwell-Specific Features
@@
 Thus developers are required to adopt entirely new instruction sequences (tcgen05.ld, tcgen05.st, tcgen05.cp).
-Second, the performance implications of this new memory tier remain uncharacterized, leaving application developers without guidance on when and how to leverage TMEM effectively.
+Second, the performance implications of this new memory tier remain characterized, leaving application developers without guidance on when and how to leverage TMEM effectively.
 Our work addresses this gap by providing the first comprehensive empirical analysis of TMEM performance characteristics and their impact on real-world computational kernels.
 Our methodology comprises of three evaluation strategies.
-Establish performance baselines by comparing memory access latencies between traditional shared memory and TMEM using pointer-chase benchmarks.
-This microbenchmark isolates latency effects by creating dependent memory accesses that prevent pipeline overlap, revealing the fundamental access costs of each memory tier.
+Establish performance baselines by comparing memory access latencies between traditional shared memory and TMEM, using pointer-chase benchmarks [28].
+To isolates access latency of each memory tier, we use dependent pointer-chase loads by creating dependent memory accesses that prevent pipeline overlap.
 Systematically compare the new TMEM data movement instructions (tcgen05.* family) against their predecessors across varying access patterns.
@@
 Latency captures complete decompression time including memory transfers.
+To isolate DE behavior from compression overhead, all datasets are pre-compressed on the CPU and the benchmark measures only device-side decompression.
 Each measurement averages 1000 iterations after 100-iteration warmup to ensure thermal and cache stability.
 We generate synthetic datasets with varying entropy: random data (incompressible, 1.00x ratio), mixed alphanumeric (1.98x), repetitive patterns (15.02x), and zero-filled buffers (245.45x).
-All experiments use LZ4 format to isolate compression-ratio effects from format-specific optimizations.
-Data is pre-compressed on CPU using reference implementations, then loaded to GPU memory for DE processing.
-This design separates compression algorithm effects from hardware decompression characteristics.
 We systematically vary chunk sizes (32KB, 64KB, 128KB, 256KB) and batch concurrency (1–1024 concurrent operations) to identify optimal parallelism levels.
@@
 We conduct latency and throughput measurements across varying instruction types, matrix tile shapes, and operand layouts to characterize execution pipeline behavior.
+To isolate instruction latency, we use dependency chains carried through the accumulator so each MMA depends on the prior result, preventing overlap; throughput is measured with independent MMAs to saturate the tensor-core pipeline.
 Power efficiency analysis compares compute throughput against board-level power consumption to identify energy-optimal operating points for different precision modes and tile configurations.
-To test the single thread instruction improvement instruction we issue multiple mma tile shapes all with FP16 inputs and outputs.
-We first isolate the instructions and show single instruction latency (SI-LAT).
-As well as compare against various levels of warps, which will account for wgmma only being issued at warp-group level which is 4 asynchronous warps per group.
-As well as using the TMEM as minimal as allowed, which means only for the accumulator, denoted as SS.
 IV-A4 Extended Precision Characterization
-Unlike prior work focusing on FP8, FP16, and INT8 tensor operations [Fasi2021, Markidis2018], we develop the first systematic benchmarks for Blackwell’s FP4 and FP6 MMA instructions using the tcgen05 PTX opcode with e2m1 (FP4), e3m2 (FP6), and e2m3 (FP6) encoding formats.
-Our key methodological contribution is a dependency-chain approach that isolates true instruction latency for these ultra-low-precision operations.
+Unlike prior work focusing on FP8, FP16, and INT8 tensor operations [6, 18], we develop the first systematic benchmarks for Blackwell’s FP4 and FP6 MMA instructions using the tcgen05 PTX opcode with e2m1 (FP4), e3m2 (FP6), and e2m3 (FP6) encoding formats.
+We use dependency chains carried through the destination operand to prevent independent issues and expose true dependent latency for each FP4/FP6 instruction variant.
 IV-A5 Workflow Benchmarks
 To assess each of these individual features as well as the whole B200 we develop integrated workloads that exercise multiple architectural innovations simultaneously.
-First, we select the Mistral model family [jiang2023mistral7b] as an LLM for several reasons: (1) Mistral-7B provides a representative dense decoder architecture with performance comparable to larger models, (2) Mixtral-8x7B’s Mixture-of-Experts (MoE) architecture exercises different dataflow patterns that stress Blackwell’s memory hierarchy, and (3) the Mistral family’s public availability enables reproducibility.
+First, we select the Mistral model family [12] as an LLM for several reasons: (1) Mistral-7B provides a representative dense decoder architecture with performance comparable to larger models, (2) Mixtral-8x7B’s Mixture-of-Experts (MoE) architecture exercises different dataflow patterns that stress Blackwell’s memory hierarchy, and (3) the Mistral family’s public availability enables reproducibility.
 The architectural diversity spanning dense (Mistral-7B) to sparse MoE (Mixtral-8x7B, Mixtral-8x22B) provides comprehensive coverage of modern LLM deployment scenarios.
 Next, we develop custom matrix multiplication kernels using FP64 to measure the realistic performance for scientific workloads.
-In addition, we run STREAM Triad [McCalpin1995] for memory bandwidth and SpMV tests with real-world data to benchmark the DE.
-Finally, we measure end-to-end training performance using mixed-precision training with ResNet-50 [he2015deepresiduallearningimage], and GPT-1.3B [gao2020pile].
+In addition, we run STREAM Triad [20] for memory bandwidth and SpMV tests with real-world data to benchmark the DE.
+Finally, we measure end-to-end training performance using mixed-precision training with ResNet-50 [8], and GPT-1.3B [7].
 Our PTX-microbenchmark methodology, detailed above, provides empirical performance data unavailable in existing simulation frameworks of the B200’s features.
@@
 In this Section, we present a comparative evaluation of the memory subsystems, including TMEM and DE, through microbenchmarking methodologies that measure latency, saturation behavior, and sensitivity to access patterns.
-These new characteristics could fundamentally alter data movement patterns and memory bandwidth utilization compared to prior architectures.
+These new characteristics can significantly change data movement patterns and effective bandwidth utilization for workloads that (i) stage tensor operands/accumulators through TMEM and (ii) stream compressed inputs through the DE, compared to prior architectures.
 V-A Tensor Memory (TMEM)
 The TMEM is a dedicated 256KB on-chip memory per SM designed solely for Tensor Core operations.
-Structured as a 2D array of 512 columns by 128 lanes of 32-bit cells, TMEM uses a lane-column addressing scheme [NVIDIA2024_Blackwell].
-The TMEM separates tensor core storage from registers, enabling intermediate matrix results to persist across warp groups and while reducing reliance on global or shared memory.
-Our latency characterization reveals that TMEM achieves 420 clock cycles for end-to-end memory access in cache-miss scenarios, representing a 58% reduction compared to Hopper’s 1000-cycle global memory latency.
-This improvement stems from TMEM’s dedicated arbitration logic that bypasses the L2 cache partitioning contention inherent in traditional memory hierarchies.
-More critically, TMEM provides 16 TB/s read bandwidth and 8 TB/s write bandwidth per SM, and this bandwidth operates additively with L1/SMEM bandwidth rather than competing for the same resources.
-In our microbenchmarks using back-to-back MMA operations on FP8 data, TMEM sustained 8 TB/s memory bandwidth matching HBM3e peak performance, achieving a 2.1× improvement over conventional ld.global paths which plateau at 3.8 TB/s due to L1/L2 traversal overhead.
+Structured as a 2D array of 512 columns by 128 lanes of 32-bit cells, TMEM uses a lane-column addressing scheme [22].
+The TMEM separates tensor core storage from registers, enabling intermediate matrix results to persist across warp groups and reducing reliance on registers or shared memory.
 The transition to TMEM necessitates entirely new instruction sequences, as traditional data movement instructions cannot interface with this memory tier.
-On Hopper, the standard pipeline for tensor operations relied on cp.async.bulk.tensor.2d for asynchronous 2D tile copies from global to shared memory, followed by ldmatrix or wmma.load to stage operands into registers before MMA execution.
-These instructions formed a well-understood pipeline: global memory copy engines populate shared memory tiles, barrier synchronization ensures data availability, and explicit load instructions would transfer operands to the register file.
+On Hopper, the standard pipeline used cp.async.bulk.tensor.2d (or TMA) for global-to-shared tile copies, then ldmatrix or wmma.load to stage operands; the wgmma instruction accepts operand A from registers or shared memory and operand B from shared memory (B is not staged into registers before MMA).
 On Blackwell, the tcgen05 instruction family replaces this entire sequence.
@@
 Our instruction-level analysis across varying operand sizes and access strides reveals critical performance characteristics.
-TMEM achieves optimal efficiency at 64×64 element tiles (4KB for FP8 precision), aligning with the 256KB SMEM capacity per SM and fully utilizing the 1024-bit memory interface width.
-This represents a significant departure from Hopper’s 32×32 optimal tile size, requiring algorithmic adjustments for kernels transitioning to Blackwell.
-Tiles smaller than 32×32 elements underutilize the wide memory interface, achieving only 45% of peak bandwidth, while tiles larger than 128×128 elements trigger multi-phase transfers that introduce pipeline stalls and reduce effective throughput by 30%.
-These bandwidth saturation points directly inform optimal kernel design: matrix multiplication kernels should decompose computations into 64×64 tiles to maximize TMEM utilization, and chained operations such as those in transformer attention mechanisms ( Q×KT followed by softmax and value multiplication) should maintain intermediate results in TMEM to exploit the 16 TB/s read bandwidth for subsequent operations.
-Traditional Hopper operations exhibits a serial dependency chain: global memory fetch, L2 traversal, shared memory write, barrier wait, register load, and finally MMA execution.
-Each stage introduced both latency and bandwidth competition.
-Blackwell’s TMEM operates with independent address generation units that pre-fetch tensor tiles directly into staging buffers, enabling the tcgen05.mma instructions to overlap data movement with computation.
-While a warp group executes tensor operations on one TMEM tile, the copy engines asynchronously populate the next tile, achieving near-perfect double buffering.
-This pipeline efficiency becomes particularly evident in workloads with producer-consumer relationships between MMA operations.
-For chained matrix multiplications where , keeping the intermediate result in TMEM eliminates an estimated 12 TB of data movement per second on a fully utilized SM compared to Hopper’s approach of writing back to global memory.
-Power efficiency measurements reveal nuanced trade-offs in TMEM utilization.
-For kernels that stage Matrix-D accumulators in TMEM versus traditional shared memory, we observe a 15% reduction in board-level power consumption at equivalent compute throughput for large matrix dimensions (2048×2048).
-This efficiency gain results from reduced L2 cache thrashing and lower DRAM traffic as intermediate results remain on-chip.
-However, for smaller problem sizes where the entire working set fits in L1 cache, forcing TMEM allocation introduces marginal overhead from the additional copy operations, resulting in a 3-5% power increase.
-These measurements establish clear guidelines: TMEM should be prioritized for multi-stage tensor pipelines with large working sets, while traditional memory hierarchies remain optimal for single-shot operations on small matrices.
+TMEM achieves optimal efficiency at 6464 element tiles (4KB for FP8 precision), aligning with the 256KB TMEM capacity per SM and fully utilizing the 1024-bit memory interface width.
+(PTX wgmma matrix shapes require M64, e.g.
+m64n8k16 through m64n256k16.) Tiles smaller than 3232 elements underutilize the wide memory interface in our TMEM microbenchmarks, while tiles larger than 128128 elements can trigger multi-phase transfers.
+These bandwidth saturation points directly inform optimal kernel design: matrix multiplication kernels should decompose computations into 64×64 tiles to maximize TMEM utilization, and chained operations such as those in transformer attention mechanisms ( followed by softmax and value multiplication) should maintain intermediate results in TMEM to exploit the 16 TB/s read bandwidth for subsequent operations.
+Traditional Hopper operations exhibit a serial dependency chain: global memory fetch (e.g.
+TMA or cp.async.bulk), shared memory write, barrier wait, then wgmma consuming operands from shared (and optionally registers).
+On Blackwell, TMEM is software-managed via tcgen05.alloc, tcgen05.cp, tcgen05.ld/tcgen05.st; software uses tcgen05.cp to copy data into TMEM from shared memory, and tcgen05.mma can then overlap with subsequent tcgen05.cp for double-buffered pipelines.
+For chained matrix multiplications where , keeping the intermediate result in TMEM eliminates a substantial amount of off-chip traffic relative to a Hopper-style design that writes intermediates back to global memory.
+Under bandwidth and traffic assumptions, this corresponds to an estimated of data movement per SM that can be avoided on a fully utilized Blackwell SM.
 V-B Decompression Engine (DE)
@@
 This indicates the dedicated optimization path for the DE.
-Most notably, Bitcomp achieves exceptional output throughput of 462.4 GB/s with minimal latency of 0.227ms, likely benefiting from integer-specific optimizations tailored for scientific workloads involving numerical data.
+Most notably, Bitcomp achieves exceptional output throughput of 462.4 GB/s with minimal latency of 0.227ms in our measurements, likely benefiting from integer-specific optimizations tailored for scientific workloads involving numerical data.
 All tested formats achieve sub-millisecond decompression latency ranging from 0.227 to 1.251ms.
@@
 | Zeros | 245.45x | 0.85 | 209.83 | 0.500 |
-The data in Table II demonstrates that the DE is fundamentally bottlenecked by compressed input bandwidth rather than decompression compute capacity.
-Random data with no compression (1.00x ratio) achieves peak input throughput of 173.23 GB/s, while highly compressed zeros data with 245.45x compression ratio drops to merely 0.85 GB/s input throughput.
-This inverse relationship indicates that the hardware must process significantly more complex decompression operations for highly compressed data, consuming proportionally more cycles per input byte.
-Despite the dramatic variation in input processing rates, output throughput remains remarkably stable at approximately 160-220 GB/s across all data patterns, with peak output performance of 219.80 GB/s achieved on repetitive data.
-This consistency suggests the presence of an internal decompression throughput ceiling of approximately 200 GB/s, indicating that the DE architecture prioritizes sustained output bandwidth over input processing rate maximization.
-The stable output performance across varying compression ratios demonstrates a decompression-throughput bounded design rather than an input-bandwidth bounded architecture.
-The latency characteristics shown in Table II remain consistently low across all data patterns (0.477-0.660ms for 100MB datasets), demonstrating that the hardware maintains predictable response times regardless of compression complexity.
+The data in Table II indicate that, for LZ4 on B200, decompressed output bandwidth is the dominant limiter rather than compressed-input bandwidth or DE compute capacity.
+Across patterns, the DE sustains roughly 170–220 GB/s of decompressed output while input bandwidth varies with compression ratio.
+For incompressible data (1.00), the DE effectively behaves as a pass-through, so compressed input throughput closely matches the output rate (173.23 GB/s).
+For highly compressible data, each compressed byte expands into many decompressed bytes; because output bandwidth remains near its limit, the compressed input rate necessarily decreases (0.85 GB/s at 245), following an approximate trend where is the compression ratio.
+Thus incompressible data achieves higher input throughput not because input bandwidth is the bottleneck, but because no expansion occurs and the DE can forward bytes at its maximum observed output rate.
+Despite large variation in input processing rates, decompressed output throughput stays within approximately 160–220 GB/s across all patterns, with peak output performance of 219.80 GB/s on repetitive data.
+This consistency suggests that, under our test conditions, DE performance is primarily bounded by available output bandwidth rather than by the cost of decompression itself.
+The stable output performance across widely varying compression ratios is consistent with a decompression-throughput–bounded design for these LZ4 workloads.
+The latency characteristics shown in Table II remain consistently low across all data patterns (0.477–0.660ms for 100MB datasets), demonstrating that the hardware maintains predictable response times regardless of compression complexity.
 This temporal consistency is crucial for real-time applications where predictable performance is more important than peak throughput, and suggests that the DE implements sophisticated workload balancing mechanisms to maintain consistent service levels.
@@
 | (GB/s) | (Concurrent Ops) | (Batch Size) | vs Sequential | |
-| 32 KB | 53.8 | 16 | 1024 | 71.95x |
-| 64 KB | 85.7 | 1 | 1024 | 69.81x |
-| 128 KB | 118.7 | 8 | 256 | 41.88x |
-| 256 KB | 151.6 | 4 | 1024 | 47.20x |
-The pipeline depth analysis in Table III reveals a clear inverse relationship between chunk size and optimal concurrency levels, with pipeline depth decreasing from 16 concurrent operations for 32KB chunks to 4 concurrent operations for 256KB chunks.
-This pattern indicates that larger chunks consume proportionally more hardware resources or memory bandwidth per operation, forcing the system to reduce parallelism to maintain optimal efficiency.
-The consistent efficiency threshold of 85% across different chunk sizes suggests that the DE implements sophisticated resource management to prevent performance degradation due to resource contention.
-Peak throughput scaling demonstrates the DE’s exceptional parallel processing capabilities, with aggregate throughput increasing from 53.8 GB/s for small chunks to 151.6 GB/s for large chunks.
-This scaling behavior indicates that while optimal concurrency decreases with chunk size, the increased data processing per operation more than compensates for reduced parallelism.
-The maximum speedups of 40-70x over sequential processing demonstrate true hardware-level parallelism, enabling the simultaneous processing of dozens to hundreds of decompression streams depending on workload characteristics.
-The saturation point analysis reveals different scaling behaviors across chunk sizes, with most configurations continuing to show throughput improvements up to 1024 concurrent operations, albeit with significantly reduced efficiency beyond the pipeline depth threshold.
-For 128KB chunks, saturation occurs earlier at 256 concurrent operations, suggesting that larger chunks create memory bandwidth or resource pressure that prevents effective scaling to higher concurrency levels.
-This behavior provides clear guidance for application developers on optimal batching strategies to maximize hardware utilization while maintaining high efficiency.
-Single-request performance varies significantly with chunk size, ranging from 0.75 GB/s for small chunks to 3.21 GB/s for large chunks, demonstrating that the DE’s baseline performance scales with data granularity.
-However, the true power of the hardware becomes apparent through batching, where aggregate throughput reaches up to 151.6 GB/s for 256KB chunks processed concurrently.
-This scaling behavior indicates that applications must carefully balance chunk size against concurrency levels to achieve optimal performance.
-Hardware efficiency remains above 90% up to the pipeline depth threshold for each chunk size, providing clear guidance for optimal operating points.
-Beyond the pipeline depth, efficiency degrades as concurrency exceeds optimal hardware utilization, with large batches showing continued throughput increases but at drastically reduced efficiency.
-This efficiency profile suggests that applications should target the discovered pipeline depth for their specific chunk sizes to maximize both throughput and resource utilization efficiency.
-The relationship between chunk size, pipeline depth, and peak throughput reveals fundamental memory bandwidth limitations within the DE architecture.
-As chunk sizes increase, the reduced pipeline depth combined with higher per-operation throughput suggests that the hardware is primarily memory bandwidth limited rather than compute limited.
-This architectural characteristic explains why the DE prioritizes sustained decompression throughput over input processing rate, as consistent memory bandwidth utilization becomes the primary performance determinant.
-Based on our empirical characterization, the optimal utilization strategy varies significantly with application requirements and data characteristics.
-For applications processing numerous small files, utilizing 32KB chunks with 16 concurrent operations maximizes aggregate bandwidth while maintaining high efficiency (90%).
-This configuration achieves 53.8 GB/s aggregate throughput with minimal per-operation latency, making it ideal for real-time data ingestion pipelines.
-Large file processing applications should leverage 256KB chunks with 4 concurrent operations to achieve maximum per-operation throughput of 151.6 GB/s.
-While this configuration supports fewer concurrent streams, the higher per-operation bandwidth more than compensates for reduced parallelism, making it optimal for applications processing large datasets or files.
-Scientific computing workloads can further optimize performance by utilizing the Bitcomp format, which achieves exceptional 462.4 GB/s output throughput with minimal latency (0.227ms) for numerical data processing.
-These findings enable developers to optimize applications for maximum DE utilization by selecting appropriate formats, chunk sizes, and concurrency levels based on specific workload requirements.
-The transformation of previously CPU-bound decompression operations into GPU-accelerated pipelines capable of sustaining 100+ GB/s throughput fundamentally changes the economics of data-intensive computing, enabling real-time processing of compressed data streams and dramatically reducing time-to-insight for applications across AI, HPC, and analytics domains.
-The B200 DE thus establishes a new performance baseline for hardware-accelerated data processing that will likely influence future architectural developments in specialized computing accelerators.
+| 32 KB | 55.84 | 1 | 1024 | 89.11x |
+| 64 KB | 71.70 | 2 | 512 | 60.61x |
+| 128 KB | 87.67 | 8 | 256 | 35.77x |
+| 256 KB | 112.10 | 8 | 256 | 47.19x |
+The pipeline depth analysis in Table III shows that, on B200 with our nvCOMP-based benchmark, small and medium chunks (32–64 KB) reach near-peak throughput with shallow pipeline depths (1–2), while larger chunks (128–256 KB) benefit from deeper concurrency (depth 8).
+Rather than a strictly monotonic “smaller chunk higher optimal depth” pattern, these results indicate that the effective concurrency needed to saturate the DE depends on both chunk size and the implementation’s internal scheduling strategy.
+Peak throughput increases from 55.84 GB/s at 32 KB to 112.10 GB/s at 256 KB, even though the optimal depth grows from 1–2 for 32–64 KB to 8 for 128–256 KB.
+This suggests that, for small chunks, the DE and memory system can saturate with only a few in-flight operations per stream, whereas larger chunks require additional parallelism to hide per-request overheads and fully utilize available bandwidth.
+Maximum observed speedups relative to a single sequential request range from roughly 35.8 to 89.1, confirming substantial benefit from batching and concurrency.
+The saturation-point column further illustrates how batching behavior changes with chunk size.
+For 32 KB, throughput continues improving up to a batch size of 1024, whereas for 128–256 KB the best efficiency is reached by 256 in-flight requests.
+Beyond these points, additional concurrency yields diminishing returns, consistent with DRAM and internal-buffer pressure limiting further scaling.
+Single-request throughput also scales with chunk size (from 0.63 GB/s at 32 KB to 3.13 GB/s at 256 KB in our runs), but aggregate throughput is dominated by how many such requests can be overlapped in hardware.
+Together, these results provide practical guidance for B200: applications with many small chunks can use low pipeline depths (1–2) and large batches, while large-chunk workloads should increase depth up to the point where throughput saturates (around depth 8 and batch size 256 in our configuration) without oversubscribing the memory subsystem.
+The combination of rising per-request throughput with chunk size and modest gains beyond the saturation point suggests that, at high utilization, DE performance is primarily limited by downstream memory bandwidth rather than by decompression compute.
+Based on our empirical characterization, the optimal utilization strategy varies with application requirements and data characteristics.
+For applications processing numerous small files, our B200 measurements indicate that 32–64 KB chunks with shallow pipeline depth (1–2) and large batch sizes (up to roughly 512–1024 requests) are sufficient to reach near-peak aggregate bandwidth while keeping per-request latency low.
+Large-file workloads should leverage 128–256 KB chunks with deeper concurrency (depth 8) and batches around 256 requests to achieve the highest observed throughput.
 VI GPU Cores Microarchitecture
@@
 Below, we describe in detail the microarchitecture of the tensor core specifically the tcgen05 instructions, CTA pair scheduling, and the extended precision support.
-VI-A Fifth-Generation Tensor Cores
+VI-A 5th-Generation Tensor Cores
 Previous studies show that tensor core PTX instruction compiles to a set of SASS instructions (HMMA, HGMMA, QGMMA, IGMMA, or BGMMA) depending on operand precision.
@@
 This represents a departure from Hopper’s unified wgmma approach, enabling precision-specific optimizations at the hardware level.
-| Precision | tcgen05.mma | wgmma |
-| FP64 | DMMA | DMMA |
-| FP32 | HMMA | HGMMA |
-| FP4 | OMMA | N/A |
-| FP8 | QMMA | QGMMA |
-| INT4, INT8 | IMMA | IGMMA |
-With the shift from warp-group-level (wgmma, 128 threads) on Hopper to warp-level (tcgen05.mma, 32 threads) execution. Our benchmarks reveal in Table V the latency implications of this design choice.
+| Precision | PTX (tcgen05.mma) | SASS | wgmma |
+| FP16, BF16 | kind::f16 | HMMA | HGMMA |
+| FP32, TF32 | kind::tf32 | HMMA | HGMMA |
+| FP8 | kind::mxf8 / f8f6f4 | QMMA | QGMMA |
+| FP6 | kind::mxf6 | QMMA | N/A |
+| FP4 | kind::mxf4 / mxf4nvf4 | OMMA | N/A |
+| INT4, INT8 | kind::i8 | IMMA | IGMMA |
+| FP64 | not supported | —∗ | DMMA |
+With the shift from warp-group-level (wgmma, 128 threads) on Hopper to warp-level (tcgen05.mma, 32 threads) execution, our benchmarks reveal in Table V the latency implications of this design choice.
+SI-LAT values are from dependency-chain microbenchmarks (accumulator-carried dependency).
 | Instruction | Tile Shape | Scope | SI-LAT (cycles) |
@@
 | tcgen05.mma | m256n256k16 | Warp | 11.4 |
-Our measurements show Blackwell achieves 2.9–11.6 lower single-instruction latency than Hopper.
-Crucially, this latency remains nearly constant across tile sizes (11.0–11.4 cycles), whereas Hopper scales linearly with tile width.
-This confirms Blackwell implements a different pipeline architecture where tile size affects throughput but not latency—indicative of a spatial array design rather than Hopper’s temporal pipelining.
-In addition, the warp-level granularity enables finer-grained scheduling and reduced synchronization overhead.
-In Hopper, four warps must synchronize for each wgmma operation; Blackwell eliminates this requirement.
-Our profiling shows this reduces scheduler stalls by 18–23% in memory-bound kernels where Tensor Core utilization is limited by data availability rather than compute capacity.
+The data in Table V are consistent with Blackwell achieving 2.9–11.2 lower single-instruction latency than Hopper for the tile sizes measured.
+Latency remains nearly constant across tile sizes (11.0–11.4 cycles), whereas Hopper scales linearly with tile width.
+The reported cycle counts reflect instruction-level latency from dependency-chain microbenchmarks.
+We hypothesize that tile size affects throughput rather than latency; pipeline depth is not directly measured, and the constant latency across tiles is consistent with a spatial array design rather than Hopper-style temporal pipelining.
+In addition, the warp-level granularity enables finer-grained scheduling and reduced synchronization overhead: in Hopper, four warps must synchronize for each wgmma operation, whereas Blackwell eliminates this requirement.
+In memory-bound kernels where tensor core utilization is limited by data availability, reduced warp-level synchronization can improve scheduler efficiency; quantitative stall percentages are environment- and workload-dependent.
+Class C warp-group scaling benchmarks (Hopper: warp-groups per SM for wgmma; B200: warps per CTA for tcgen05.mma) collect Nsight Compute scheduler-stall and tensor-utilization counters.
 | Input (A/B) | Accum (C/D) | Shape | Latency | Throughput |
@@
 | INT8 | INT32 | m64n8k16 | 11.9 | 3928.5 |
-Expanding our analysis across supported precisions, our results in Table VI shows that despite 177 throughput difference between FP64 (44.8 TFLOPS) and FP4 (7702.5 TFLOPS), latency varies by only 1.27 (11.2–14.2 cycles).
-This confirms throughput scaling is achieved through increased parallelism (wider datapaths) rather than deeper pipelining.
-Hence, Blackwell prioritizes consistent low-latency operation across all precisions, enabling predictable performance regardless of quantization level.
+Expanding our analysis across supported precisions, our results in Table VI show that across the measured precisions, throughput spans 8.2 (481.6 to 3928.5 TFLOPS or TOPS), while latency varies by only 1.12 (11.2–12.6 cycles).
+This is consistent with throughput scaling being achieved through increased parallelism (wider datapaths) rather than deeper pipelining.
+Hence, the data suggest Blackwell prioritizes consistent low-latency operation across all precisions, enabling predictable performance regardless of quantization level.
 Comparing FP16 inputs with FP16 vs.
-FP32 accumulators reveals a critical bottleneck that FP32 accumulation halves throughput (1929.2 → 964.6 TFLOPS).
+FP32 accumulators reveals a critical bottleneck that FP32 accumulation halves throughput (964.8 → 482.4 TFLOPS).
 This indicates the accumulator datapath, not the multiply units, limits throughput, also noted in the previous sections.
 Meaning a trade-off is that applications requiring high numerical precision must sacrifice 50% performance, while inference workloads using FP16 accumulators achieve maximum throughput.
-INT8 (3927.1 TOPS) slightly exceeds FP8 (3851.4 TFLOPS), while FP4 (7702.5 TFLOPS) outperforms FP8.
+INT8 (3928.5 TOPS) exceeds FP8 (1912.8 TFLOPS), and FP4 (3850.1 TFLOPS) outperforms FP8.
 This advantage suggests both integer and floating-point operations share the same execution units, with integer formats requiring marginally simpler control logic.
-With most precisions showing a similar latency this confirms that the pipeline is similar and increased parallelism improves throughput.
+With most precisions showing a similar latency (Table VI), this is consistent with a similar pipeline and increased parallelism improving throughput.
 VI-B Extended Precision Support: FP4 and FP6
 One of Blackwell’s most significant improvements is native hardware support for FP4 and FP6 (6-bit floating-point) data types.
-Table VI shows the different supported precision formats of the Blackwell Tensor Cores.
+FP4 tensor operations are verified via CUTLASS disassembly, which shows the OMMA instruction.
+FP6 is documented to map kind::mxf6 to QMMA per the PTX ISA (Table IV).
+Table VI shows latency (12.6 cycles for FP4/FP16, 12.3 for FP6/FP16) and throughput (3850.1 and 2567.2 TFLOPS at m64n8k16).
+The FP4/FP16 and FP6/FP16 configurations use FP32 accumulation (PTX .f32 output); reported throughput values include accumulator datapath costs.
 The FP4 format uses 1 sign bit, 2 exponent bits, and 1 mantissa bit as e2m1.
-Available on Blackwell are MXFP4, microscaling floating-point [Rouhani2023MicroscalingDF], and NVFP4, from NVIDIA.
+Available on Blackwell are MXFP4, microscaling floating-point [24], and NVFP4, from NVIDIA.
 MXFP4 enhances low-precision training by dividing data into blocks of size 32, with each using a scale with E8M0 format.
@@
 Work by Chmiel et.
-al [chmiel2025fp4wayfullyquantized] provides a more indepth comparison.
-While this extremely limited precision might seem impractical, recent quantization research has demonstrated that for inference workloads FP4 can maintain acceptable accuracy [dettmers2023case4bitprecisionkbit].
-Blackwell’s FP4 support includes hardware dequantization logic that converts FP4 values to higher precision (typically FP8 or FP16) during matrix multiplication.
-This allows FP4 storage and bandwidth savings while maintaining computational precision.
+al [1] provides a more indepth comparison.
+While this extremely limited precision might seem impractical, recent quantization research has demonstrated that for inference workloads FP4 can maintain acceptable accuracy [4].
+Blackwell’s FP4 path integrates any necessary precision conversion within the MMA pipeline; reported latency (12.6 cycles) and throughput (96.2% of peak) include the complete FP4 execution path.
+We do not measure dequantization as a standalone operation.
 On the other hand, FP6 provides a middle ground, using 1 sign bit, 3 exponent bits, and 2 mantissa bits.
 This format offers significantly better dynamic range than FP4 while still providing 1.33 memory and bandwidth savings compared to FP8.
-Table VII shows effective throughput for different precision modes.
+Table VII shows effective throughput for different precision modes (FP6: 5134.4 TFLOPS, 96.0% peak; FP4: 7700.2 TFLOPS, 96.2% peak).
 | Precision | B200 | % Peak | H200 | Speedup |
 | FP64 | 44.8 | 99.6% | 34.0 | 1.32 |
-| FP32 | 481.2 | 96.2% | 378.4 | 1.27 |
+| FP32 | 482.0 | 96.4% | 378.4 | 1.27 |
 | TF32 | 964.5 | 96.5% | 756.9 | 1.27 |
-| BF16 | 1926.8 | 96.3% | 1513.5 | 1.27 |
-| FP16 | 1929.2 | 96.5% | 1515.2 | 1.27 |
-| FP8 | 3851.4 | 96.3% | 3026.9 | 1.27 |
-| FP6 | 5134.8 | 95.8% | N/A | New |
-| FP4 | 7702.5 | 96.3% | N/A | New |
-| INT8 | 3927.1 | 98.2% | 3088.4 | 1.27 |
-Our measurements confirm architectural specifications.
-At matrix dimensions, Blackwell achieves 3851.4 TFLOPS in FP8 mode, representing 96.3% of theoretical peak.
-For FP4 operations, achieved throughput of 7702 TFLOPS represents 96.3% of peak.
-With 96-99% of theoretical peak across all precisions this shows Tensor Cores are not the bottleneck, where memory bandwidth and kernel launch overhead is.
+| BF16 | 1926.4 | 96.3% | 1513.5 | 1.27 |
+| FP16 | 1929.6 | 96.5% | 1515.2 | 1.27 |
+| FP8 | 3850.6 | 96.3% | 3026.9 | 1.27 |
+| FP6 | 5134.4 | 96.0% | N/A | New |
+| FP4 | 7700.2 | 96.2% | N/A | New |
+| INT8 | 3928.5 | 98.2% | 3088.4 | 1.27 |
+Our achieved throughput is consistent with architectural specifications when compared to theoretical peak (peak ).
+At the reported matrix dimensions, we observe 3850.6 TFLOPS in FP8 mode (96.3% of theoretical peak) and 7700.2 TFLOPS in FP4 (96.2% of peak), see Table VII.
+With 96–99% of theoretical peak across all precisions in our runs, tensor cores are not the bottleneck; memory bandwidth and kernel launch overhead dominate.
 In Section VII, we analyze the use of tensor cores, TMEM, and DE for different real-world workloads.
 | Model | Precision | B200 tok/s | H200 tok/s | Speedup | B200 BW% | H200 BW% | Perplexity | PPL |
-| Mistral-7B | FP16 | 45,200 | 28,500 | 1.59 | 67.3 | 71.2 | 6.82 | — |
-| FP8 | 78,400 | 49,200 | 1.59 | 58.4 | 62.8 | 6.95 | +1.9% | |
+| Mistral-7B | FP16 | 56,028 | 28,500 | 1.97 | 67.3 | 71.2 | 6.82 | — |
+| FP8 | 57,125 | 49,200 | 1.16 | 58.4 | 62.8 | 6.95 | +1.9% | |
 | FP4 | 112,800 | N/A | N/A | 47.6 | N/A | 7.38 | +8.2% | |
-| Mixtral-8x7B | FP16 | 28,600 | 18,100 | 1.58 | 72.1 | 76.4 | 5.94 | — |
-| FP8 | 51,200 | 32,400 | 1.58 | 61.8 | 65.2 | 6.08 | +2.4% | |
+| Mixtral-8x7B | FP16 | 31,033 | 18,100 | 1.71 | 72.1 | 76.4 | 5.94 | — |
+| FP8 | 51,200† | 32,400 | 1.58 | 61.8 | 65.2 | 6.08 | +2.4% | |
 | FP4 | 76,900 | N/A | N/A | 49.1 | N/A | 6.48 | +9.1% | |
@@
 Our analysis quantifies the performance benefits of architectural innovations in the NVIDIA B200 compared to the H200 baseline.
+The reported speedups are directly informed by our microbenchmark findings: TMEM latency and bandwidth (Section V) explain improved fused-attention and tensor-resident accumulation; DE output-bandwidth behavior (Section V) underlies SpMV and compressed-data workloads; and 5th-generation tensor core latency and precision scaling (Section VI) drive the LLM and training gains below.
 VII-A Experimental Methodology
@@
 Energy consumption is monitored using the NVIDIA Management Library (NVML) API with 10ms sampling intervals to provide high-resolution power profiling.
+Software and environment.
+All experiments use CUDA 12.6 (B200 and H200), cuBLAS/cuBLASLt from the same toolkit, and PyTorch 2.4 with Transformer Engine where applicable for LLM inference and training.
+DGEMM and STREAM runs use the same CUDA driver (560.x) and compiler (nvcc 12.6).
+We report L2 cache hit rates (e.g., in Section VII.B) from NVIDIA Nsight Compute memory-workload analysis; the reported 68–84% range is derived from the l2tex_throughput and global memory throughput counters over the kernel execution.
 VII-B Large Language Model Inference
@@
 Our findings reveal that lower-precision formats achieve performance gains over the FP16 baseline.
-Specifically, FP8 and FP4 deliver throughput improvements of and , respectively, for Mistral-7B.
+Specifically, FP8 and FP4 deliver throughput improvements over the FP16 baseline for Mistral-7B.
 While these gains approach the theoretical bandwidth, 2 and 4, they represent practical speedups achievable in real workloads.
@@
 Sparse mixture-of-experts models demonstrate amplified benefits from quantization compared to dense models.
-For FP4 quantization, Mixtral-8x7B achieves throughput improvement (76,900 tok/s vs 28,600 tok/s FP16 baseline) compared to for the dense Mistral-7B model.
+For FP4 quantization, Mixtral-8x7B achieves throughput improvement (76,900 tok/s vs 31,033 tok/s FP16 baseline,) compared to for the dense Mistral-7B model.
 This additional benefit stems from quantization enabling more efficient expert weight caching and reduced overhead in the expert routing mechanism.
 The B200 maintains consistent performance advantages over the H200 across all precision modes where both architectures support the format.
-For both FP16 and FP8, the B200 delivers higher throughput than the H200.
-This scaling factor reflects the combined contributions of increased SM count (), enhanced Tensor Core efficiency (), and improved effective memory bandwidth ().
+For both FP16 and FP8, the B200 delivers higher throughput than the H200 (Table VIII, Section VI).
+This scaling factor is consistent with combined contributions from increased SM count (), enhanced tensor core efficiency (; see Table VII), and improved effective memory bandwidth (); we have not directly measured this decomposition.
 Lastly, while quantization delivers substantial performance benefits, it comes with measurable but often acceptable quality degration.
@@
 | HPC DGEMM (FP64) | TFLOPS | 36.3 | 18.9 | 1.92 | Doubled FP64 units |
-| STREAM Triad | BW (TB/s) | 7.48 | 4.38 | 1.71 | HBM3e |
-| SpMV (compressed) | GFLOPS | 5.04 | 3.2 | 1.58 | Decomp engine |
-| GPT Training (1.3B) | tok/s | 14,397 | 9,240 | 1.56 | CTA pairs, TMEM, TC |
-| ResNet Training | img/s | 2,436 | 1,580 | 1.54 | 5th Gen TC, mem BW |
-| Energy Eff. (Training) | tok/s/W | 22.2 | 15.6 | 1.42 | Process, efficiency |
+| STREAM Triad (4–16 GB) | BW (TB/s) | 4.14 | – | – | HBM3e (B200 only) |
+| SpMV (compressed) | GFLOPS | 5.08 | 3.2 | 1.58 | Decomp engine |
+| GPT Training (1.3B) | tok/s | 14,363 | 9,240 | 1.55 | CTA pairs, TMEM, TC |
+| ResNet Training | img/s | 2,928 | 1,580 | 1.85 | 5th Gen TC, mem BW |
+| Energy Eff. (Training) | tok/s/W | 20.63 | 15.6 | 1.32 | Process, efficiency |
 | Model | Batch Size | B200 | H200 | Ratio | Time to Acc | Time to Acc | Energy |
 | Throughput | Throughput | B200 (hrs) | H200 (hrs) | Eff (B200) | |||
-| ResNet-50 | 1024 | 2,436 img/s | 1,580 img/s | 1.54 | 1.05 | 1.62 | 3.77 img/s/W |
-| GPT-1.3B | 128 | 14,397 tok/s | 9,240 tok/s | 1.56 | 5,788 | 9,020 | 22.2 tok/s/W |
-| GPT-1.3B | 64 | 14,141 tok/s | 9,070 tok/s | 1.56 | 5,893 | 9,184 | 21.8 tok/s/W |
+| ResNet-50 | 1024 | 2,928 img/s | 1,580 img/s | 1.85 | 0.87 | 1.62 | 5.09 img/s/W |
+| GPT-1.3B | 128 | 14,363 tok/s | 9,240 tok/s | 1.55 | 5,788 | 9,020 | 20.63 tok/s/W |
+| GPT-1.3B | 64 | 14,121 tok/s | 9,070 tok/s | 1.55 | 5,893 | 9,184 | 20.27 tok/s/W |
 VII-C Scientific Computing Workload
@@
 Scientific computing applications present fundamentally different computational characteristics compared to deep learning workloads, needing high-precision arithmetic, sustainable memory bandwidth, and irregular access patterns.
-We evaluate dense matrix multiplication (DGEMM) performance using dobule-precision FP arithmetic (FP64), which remains essential for scientific simulations requiring numerical accuracy.
+We evaluate dense matrix multiplication (DGEMM) performance using double-precision FP arithmetic (FP64), which remains essential for scientific simulations requiring numerical accuracy.
 Table XII presents our results across varying matrix dimensions.
-The B200 achieves 36.3 TFLOPS at large matrix size, representing 80.7% utilization of its 40 TFLOPS theoretical peak [nvidia_blackwell_b200_datasheet], compared to the H200’s 18.9 TFLOPS (55.6% of 34 TFLOPS).
-The additional 45% efficiency improvement results from TMEM-enabled accumulation that reduces L2 cache traffic and improved memory access coalescing patterns that better utilize the available memory bandwidth.
+The B200 achieves 36.3 TFLOPS at large matrix size, representing 80.7% utilization of its 40 TFLOPS theoretical peak [2], compared to the H200’s 18.9 TFLOPS (55.6% of 34 TFLOPS).
+PTX tcgen05.mma does not support FP64 (documented precisions are tf32, f16/bf16, i8/u8, and f4/f6/f8), so FP64 DGEMM uses a different execution path.
+The additional 45% efficiency improvement is attributed to doubled FP64 units and improved memory access coalescing.
 VII-C2 Sustained Memory Bandwidth
 Memory-intensive scientific applications require sustained high-bandwidth data movement capabilities.
-We employ the STREAM Triad benchmark to measure achievable memory banwidth across different working set sizes, as shown in Table XIII.
-| Array Size | B200 (TB/s) | H200 (TB/s) | B200 % | H200 % |
-| 4GB | 4.134 | 2.88 | 51.7 | 60.0 |
-| 16GB | 4.141 | 2.91 | 51.8 | 60.6 |
-| 64GB | 7.42 | 4.35 | 92.8 | 90.6 |
-| 128GB | 7.48 | 4.38 | 93.5 | 91.3 |
-The results show, smaller arrays that git within the GPU’s cache hierarchy achieve lower absolute bandwidth but higher efficiency on the H200 (60-60.6%) compared to the B200 (51.7-51.8%).
-This reflects the H200’s is more tuned for smaller working sets.
-Though, for working sets that exceed cache capacity, both architectures achieve excellent memory bandwidth utilization exceeding 90%.
-The 1.71 speedup closely matches raw bandwidth ratio (8.0/4.8), demonstrating memory-bound codes scale linearly with available bandwidth.
+We employ the STREAM Triad benchmark to measure achievable memory bandwidth.
+Table XIII reports results from our B200 runs.
+Triad uses three arrays, so 64 GB and 128 GB array sizes would require device memory 192 GB and 384 GB respectively; our evaluation node had insufficient memory for those sizes, so we report only 4–16 GB.
+| Array Size | B200 (TB/s) | B200 % peak |
+| 4 GB | 4.141 | 51.8 |
+| 16 GB | 4.140 | 51.8 |
+At 4–16 GB working sets, B200 achieves 4.14 TB/s (52% of 8 TB/s peak).
+Larger arrays would be needed to approach peak bandwidth but require devices with 192 GB memory for 64 GB and 384 GB for 128 GB in our benchmark setup.
 VII-C3 Sparse Operations
@@
 | Matrix | Sparsity | GFLOPS | Speedup | Avg Time (ms) |
-| webbase-1M | 99.99% | 5.09 | 3.16× | 39.32 |
-| circuit5M | 99.95% | 4.96 | 3.16× | 201.44 |
-| ldoor | 99.98% | 5.04 | 3.16× | 71.93 |
-The decompression engine yields consistent 3.16 speedups on sparse matrices.
-Run-length encoding (RLE) achieves 8.2 compression ratio for sparse row pointer arrays.
-The dedicated decompression hardware introduces less than 5% latency overhead while providing 35% traffic reduction in memory traffic for pointer-intensive workloads.
+| webbase-1M | 99.99% | 5.08 | 3.16× | 39.32 |
+| circuit5M | 99.95% | 4.94 | 3.16× | 201.44 |
+| ldoor | 99.98% | 5.03 | 3.16× | 71.93 |
+Using compressed sparse representations (software compression), SpMV achieves an estimated 3.16 speedup over an uncompressed baseline in our benchmarks; run-length encoding (RLE) yields approximately 8.2 compression ratio for sparse row pointer arrays.
+Hardware decompression engine (DE) characterization - throughput, latency, and pipeline depth - is presented in Section V.
 VII-D Mixed-Precision Training: End-to-End Training Performance
 We present comprehensive training benchmarks across different model architectures to assess the practical impact of architectural improvements in realistic training scenarios, as summarized in Table XI.
-Training achieves consistent 1.54–1.56 speedup, decomposing into SM count (1.09), CTA pairing (1.27), and TMEM (1.26).
-Energy efficiency improves 42% for GPT training despite 14% higher power consumption.
+Training speedup decomposes into SM count (1.09), CTA pairing (1.27), and TMEM (1.26).
+Energy efficiency improves for GPT training despite higher power consumption.
 VIII Discussion
@@
 180B) but deliver 1.5–3.9 gains.
-The 256KB TMEM per SM (10% of SM memory) achieves 61–82% hit rates, validating sizing decisions.
+The 256KB TMEM per SM (10% of SM memory) achieves 61–82% hit rates in our characterization.
+Tradeoffs and drawbacks: (1) TMEM addresses register pressure and L2 traffic but requires explicit allocation and tcgen05 data movement; kernels must be rewritten for Blackwell.
+(2) The DE is output-bandwidth-limited; on highly compressible data, compressed input throughput drops (e.g., scaling), so DE use is most beneficial when decompressed output, not input bandwidth, is the bottleneck.
+(3) PTX tcgen05.mma does not support FP64; FP64 DGEMM uses a different path (e.g., doubled FP64 units), so TMEM does not directly improve FP64 scientific kernels.
 Software Ecosystem: CUDA 13.0 provides preliminary TMEM/CTA support; framework integration ongoing.
@@
 FP4/FP6 require per-layer precision selection—8.2% perplexity degradation for FP4 represents averages; some layers tolerate FP4 while others need FP8.
+Performance guidelines.
+We consolidate actionable recommendations: (a) Use TMEM for accumulator staging and multi-stage tensor pipelines (e.g., fused attention then softmax then value) when working sets exceed L1; for single-shot small matrices, minimize TMEM round-trips.
+(b) Scale DE usage with compression ratio: expect 170–220 GB/s decompressed output; compressed input rate scales inversely with ratio.
+(c) Leverage warp-level tcgen05.mma and CTA-pair execution for GEMM and LLM inference; target 6464 tiles for TMEM.
+(d) For LLM inference, FP8 balances throughput and quality; FP4 yields higher throughput with acceptable perplexity increase when validated per model.
 Deployment: For LLM inference, B200 provides 1.8–3.9 advantages; FP4 practical for 70B models.
-Training improvements (1.54–1.56) enable 33% larger batches.
+Training improvements (1.85 ResNet-50, 1.55 GPT-1.3B) enable larger batches.
 HPC gains (1.92 FP64) competitive for scientific computing.
@@
 Our work offers insights into its architectural innovations and performance behavior.
-We quantify the impact of TMEM on matrix-heavy workloads, evaluate the throughput and optimal usage of the hardware decompression engine, and analyze 5th generation Tensor Core execution via the new tcgen05 PTX instructions.
+We quantify the impact of TMEM on matrix-heavy workloads, evaluate the throughput and optimal usage of the hardware decompression engine, and analyze 5th-generation tensor core execution via the new tcgen05 PTX instructions.
 Our study further assesses FP4 and FP6 precision trade-offs, benchmarks Blackwell across diverse workloads—including LLM inference, scientific kernels, and mixed-precision training—and distills actionable performance guidelines for developers targeting this next-generation architecture.
+Our experiments primarily exercised single-die behavior; dual-die (NV-HBI) interactions and cross-die latency remain as future work.
+Acknowledgment
+This research used resources from the NVIDIA Brev Cloud Compute and the Google Cloud Provider.
+This material is based upon work supported by the U.S.
+DOE under Contract DE-FOA-0003177, S4PST: Next Generation Science Software Technologies Project.